### 04 Jan 2024 v1

# Iyama–Solberg correspondence for exact dg categories

Abstract.
We generalize the notions of $d$-cluster tilting pair and $d$-Auslander exact dg category to $d$-precluster tilting triple and $d$-minimal Auslander–Gorenstein exact dg category. We give a bijection between equivalence classes of $d$-precluster tilting triples and equivalence classes of $d$-minimal Auslander–Gorenstein exact dg categories. Our bijection generalizes Iyama–Solberg correspondence for module categories.
Key words and phrases:
extriangulated category, $d$-minimal Auslander–Gorenstein category, exact dg category, derived dg category, subcategory stable under extensions, Iyama–Solberg correspondence
2010 Mathematics Subject Classification:
18G25, 18E30, 16E30, 16E45
1. Introduction
Let $k$ be a commutative artinian ring. An Artin $k$-algebra is a $k$-algebra which is finitely generated as a $k$-module. Let $A$ be an Artin $k$-algebra. We denote by $\mathrm{mod}A$ the category of finitely generated right $A$-modules. Let $d\geq 1$ be an integer. Recall that an $A$-module $M$ is $d$-cluster-tilting [14] if
$\mathrm{add}M$ =$=\{X\in\mathrm{mod}A\mid\mathrm{Ext}^{i}(M,X)=0\text{ for $1\leq i\leq d-1$}\}$
=$=\{Y\in\mathrm{mod}A\mid\mathrm{Ext}^{i}(Y,M)=0\text{ for $1\leq i\leq d-1$}\}.$
Iyama [13] generalized the classical Auslander correspondence [1] to the so-called Auslander–Iyama correspondence between
    Morita classes of $(A,M)$, where $A$ is an Artin $k$-algebra and $M$ a $d$-cluster tilting $A$-module, and
    Morita classes of $d$-Auslander algebras $\Gamma$, i.e. an Artin $k$-algebra $\Gamma$ such that
    (1.0.1) $\mathrm{gldim}\Gamma\leq d+1\leq\mathrm{domdim}\Gamma.$
The correspondence takes $(A,M)$ to $\Gamma=\mathrm{End}_{A}(M)$. Here $\mathrm{domdim}\Gamma$ denotes the dominant dimension of $\Gamma$ and $\mathrm{gldim}\Gamma$ denotes the global dimension of $\Gamma$. Recall that the dominant dimension of an abelian category $\mathcal{A}$ is the largest $n\in\mathbb{N}\cup\{\infty\}$ such that any projective object in $\mathcal{A}$ has an injective coresolution whose first $n$ terms are also projective. The dominant dimension of $A$ is defined as the dominant dimension of $\mathrm{mod}A$.
In [6], we extended Auslander–Iyama correspondence to the setting of exact dg categories. Let us explain this in more detail. For a pair $(\mathcal{P},\mathcal{I})$, where $\mathcal{P}$ is a connective additive dg category and $\mathcal{I}$ is an additive dg subcategory of $\mathcal{P}$, and for $n\geq 0$, we denote by $\mathcal{A}^{(n)}_{\mathcal{P},\mathcal{I}}$ the full dg subcategory of $\mathrm{pretr}(\mathcal{P})$ consisting of the objects in
$\mathcal{P}\ast\Sigma\mathcal{P}\ast\cdots\ast\Sigma^{n-1}\mathcal{P}\ast\Sigma^{n}\mathcal{P}\cap\mathrm{ker}\mathrm{Ext}^{\geq 1}(-,\mathcal{I}).$
The pair $(\mathcal{P},\mathcal{I})$ is $d$-cluster tilting if $H^{0}(\mathcal{I})$ is covariantly finite in $H^{0}(\mathcal{A}^{(d)}_{\mathcal{P},\mathcal{I}})$. Following Gorsky–Nakaoka–Palu [10], for $d\geq 0$, an exact dg category $\mathcal{A}$ is $d$-Auslander if the extriangulated category $H^{0}(\mathcal{A})$ has enough projective objects and satisfies similar homological properties as in (1.0.1). Then the equivalence classes of $d$-cluster tilting pairs $(\mathcal{P},\mathcal{I})$ correspond to the equivalence classes of connective $d$-Auslander exact dg categories. We mention that for a $d$-cluster tilting pair $(\mathcal{P},\mathcal{I})$, which consists of $\mathcal{I}=\mathrm{inj}A$ for some Artin $k$-algebra $\Lambda$ and $\mathcal{P}=\mathrm{add}M$ for some $A$-module $M$, the module $M$ is $d$-cluster tilting if and only if the corresponding $d$-Auslander exact dg category is concentrated in degree zero and abelian.
The notion of $d$-precluster tilting subcategory was introduced in [15] as a generalization of the notion of $d$-cluster tilting subcategory, cf. Definition 3.9. An $A$-module $M$ is $d$-precluster-tilting if $\mathrm{add}M$ is a $d$-precluster tilting subcategory of $\mathrm{mod}A$. The notion of $d$-minimal Auslander–Gorenstein algebra is introduced in loc. cit as a generalization of the $d$-Auslander property. An Artin $k$-algebra $\Gamma$ is $d$-minimal Auslander–Gorenstein if it satisfies
$\mathrm{id}_{\Gamma}\Gamma\leq d+1\leq\mathrm{domdim}\Gamma$
where $\mathrm{id}_{\Gamma}\Gamma$ denotes the injective dimension of $\Gamma\in\mathrm{mod}\Gamma$. Iyama–Solberg [15, Theorem 4.5] showed the following
Theorem 1.1 (Iyama–Solberg correspondence).
The map $(A,M)\mapsto\Gamma=\mathrm{End}_{A}(M)$ induces a bijective correspondence between
    Morita classes of $(A,M)$, where $A$ is an Artin $k$-algebra and $M$ is a $d$-precluster tilting $A$-module, and
    Morita classes of $d$-minimal Auslander–Gorenstein $k$-algebras $\Gamma$.
In this article, we continue the work in [6] and generalize the above Iyama–Solberg correspondence to the setting of exact dg categories. Let $d\geq 0$ be an integer. Our key notion is the following
Definition 1.2 (Definition 2.4).
Let $\mathcal{K}$ be an additive dg category and $\mathcal{I}\subseteq\mathcal{P}\subseteq\mathcal{K}$ full additive dg subcategories. The triple $(\mathcal{K},\mathcal{P},\mathcal{I})$ is $d$-precluster tilting if the following properties hold
  • (1)
    $\mathcal{P}$ is connective;
  • (2)
    $H^{0}(\mathcal{I})$ is covariantly finite in $H^{0}(\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d)})$ (cf. below);
  • (3)
    $\mathcal{K}$ is extension-closed in $\mathrm{tr}(\mathcal{K})$;
  • (4)
    $\mathrm{Hom}_{\mathrm{tr}(\mathcal{K})}(\mathcal{P},\Sigma^{\geq i}\mathcal{K})=0$ and $\mathrm{Hom}_{\mathrm{tr}(\mathcal{K})}(\mathcal{K},\Sigma^{\geq i}\mathcal{P})=0$ for any $i\geq 1$;
  • (5)
    for each object $K$ in $\mathcal{K}$, there is a triangle in $\mathrm{tr}(\mathcal{K})$
    $K^{\prime}\rightarrow P\rightarrow K\rightarrow\Sigma K^{\prime}$
    where $P\in\mathcal{P}$ and $K^{\prime}\in\mathcal{K}$.
For a triple $(\mathcal{K},\mathcal{P},\mathcal{I})$, where $\mathcal{K}$ is an additive dg category and $\mathcal{I}\subset\mathcal{P}$ are connective additive dg subcategories of $\mathcal{K}$, and for $n\geq 0$, we denote by $\mathcal{B}^{(n)}_{\mathcal{K},\mathcal{P},\mathcal{I}}$ the full dg subcategory of $\mathrm{pretr}(\mathcal{K})$ consisting of the objects in
$\mathcal{K}\ast\Sigma\mathcal{K}\ast\ldots\ast\Sigma^{n-1}\mathcal{K}\ast\Sigma^{n}\mathcal{K}\cap\mathrm{ker}^{\geq 1}\mathrm{Ext}(-,\mathcal{I}).$
Theorem 1.3 (Theorem 2.8).
There is a canonical bijection between the following:
  • (1)
    equivalence classes of connective exact dg categories $\mathcal{A}$ which are $d$-minimal Auslander–Gorenstein;
  • (2)
    equivalence classes of $d$-precluster tilting triples $(\mathcal{K},\mathcal{P},\mathcal{I})$.
The bijection from $(1)$ to $(2)$ sends $\mathcal{A}$ to the triple $(\mathcal{K},\mathcal{P},\mathcal{I})$ formed by the full dg subcategory $\mathcal{P}$ on the projectives of $\mathcal{A}$, its full dg subcategory $\mathcal{I}$ of projective-injectives and the full dg subcategory $\mathcal{K}$ of $\mathcal{D}^{b}_{dg}(\mathcal{A})$ consisting of the objects $K$ which are $(d+1)$-st syzygies of objects in $\mathcal{A}$ (in particular we have $\mathrm{Ext}^{\geq 1}(K,\mathcal{P})=0$). The inverse bijection sends $(\mathcal{K},\mathcal{P},\mathcal{I})$ to the $\tau_{\leq 0}$-truncation of the dg subcategory $\mathcal{B}^{(d+1)}_{\mathcal{K},\mathcal{P},\mathcal{I}}$ of $\mathrm{pretr}(\mathcal{K})$.
We refer to Definition 2.6 for the definition of the equivalence relations mentioned in Theorem 1.3.
We give a necessary and sufficient condition for the exact dg category $\tau_{\leq 0}\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ to be concentrated in degree zero (resp. to be abelian), cf. Proposition 3.2 (resp. Proposition 3.5). We show in Subsection 3.8 that the dg category $\mathcal{A}$ in Theorem 1.3 (1) is the module category over an Artin algebra $\Gamma$ (so in particular $\mathcal{A}$ is concentrated in degree zero) if and only if the corresponding triple $(\mathcal{K},\mathcal{P},\mathcal{I})$ satisfies $\mathcal{P}=\mathrm{add}M\subset\mathrm{mod}\Lambda$ for a $d$-precluster tilting module $M$ over some Artin algebra $\Lambda$, and $\mathcal{I}$ is the full subcategory of injective $\Lambda$-modules and $\mathcal{K}$ is the full dg subcategory of $\mathcal{D}^{b}_{dg}(\mathrm{mod}\mathcal{P})$ consisting of all $(d+1)$-st syzygies of $\mathcal{P}$-modules. In particular, our Theorem 1.3 generalizes Theorem 1.1. In Subsection 3.13, we also compare our results with the correspondence obtained by Grevstad in [11].
Conventions
In this article, we fix a commutative ring $k$. A dg $k$-category $\mathcal{A}$ is additive if $H^{0}(\mathcal{A})$ is an additive category. Additive subcategories of additive categories are assumed to be closed under direct summands. For an additive $k$-category $\mathcal{C}$, we denote by $\mathrm{Mod}\,\mathcal{C}$ the category of $k$-linear functors $\mathcal{C}^{op}\rightarrow k\mbox{-}\mathrm{Mod}\,$. For each object $C\in\mathcal{C}$, we denote by $C^{\wedge}$ the representable right $\mathcal{C}$-module $\mathrm{Hom}_{\mathcal{C}}(?,C)\in\mathrm{Mod}\,\mathcal{C}$. We denote by $\mathrm{mod}\mathcal{C}$ the category of finitely presented right $\mathcal{C}$-modules. Dually, we denote by $\mathcal{C}\mbox{-}\mathrm{Mod}\,$ (resp. $\mathcal{C}\mbox{-}\mathrm{mod}$) the category of $k$-linear functors $\mathcal{C}\rightarrow k\mbox{-}\mathrm{Mod}\,$ (resp. the category of finitely presented left $\mathcal{C}$-modules). For a subcategory $\mathcal{D}$ of $\mathcal{C}$, a morphism $f:X\rightarrow Y$ in $\mathcal{C}$ is a left $\mathcal{D}$-approximation of $X\in\mathcal{C}$ if $Y\in\mathcal{D}$ and the induced map $\mathrm{Hom}_{\mathcal{C}}(Y,Z)\rightarrow\mathrm{Hom}_{\mathcal{C}}(X,Z)$ is surjective for each $Z\in\mathcal{D}$. The full subcategory $\mathcal{D}$ is cavariantly finite if any $X\in\mathcal{C}$ has a left $\mathcal{D}$-approximation. If $\mathcal{X}$ is a class of objects in $\mathcal{C}$, we write ${{}^{\perp}\mathcal{X}}\coloneqq\{C\in\mathcal{C}\mid\mathrm{Hom}_{\mathcal{C}}(C,X)=0\text{ for any $X\in\mathcal{X}$}\}$. For a dg $k$-category $\mathcal{P}$, we denote by $\mathrm{pretr}(\mathcal{P})$ its pretriangulated hull [3, 8, 4] and by $\mathrm{tr}(\mathcal{P})=H^{0}(\mathrm{pretr}(\mathcal{P}))$ its triangulated hull. A full additive dg subcategory $\mathcal{I}\subseteq\mathcal{P}$ is covariantly finite if $H^{0}(\mathcal{I})$ is a covariantly finite subcategory of $H^{0}(\mathcal{P})$. For an exact dg category $\mathcal{A}$, we denote by $\mathcal{D}^{b}_{dg}(\mathcal{A})$ its (bounded) dg derived category and by $\mathcal{D}^{b}(\mathcal{A})=H^{0}(\mathcal{D}^{b}_{dg}(\mathcal{A}))$ its bounded derived category [5]. We denote by $\mathrm{dgcat}$ the category of small dg $k$-categories and by $\mathrm{Hqe}$ the localization of $\mathrm{dgcat}$ with respect to the quasi-equivalences [17].
Acknowledgements
The author thanks his Ph. D. supervisor Bernhard Keller and his Ph. D. co-supervisor Xiao-Wu Chen for their constant support and their encouragements. The author thanks Yann Palu, Mikhail Gorsky and Norihiro Hanihara for helpful discussions.
2. Iyama–Solberg correspondence
Throughout this section, let $d\geq 0$ be an integer.
Definition 2.1 ([10, Definition 3.5]).
Let $(\mathcal{C},\mathbb{E},\mathfrak{s})$ be an extriangulated category with enough projectives. We define its dominant dimension $\mathrm{dom.dim}(\mathcal{C},\mathbb{E},\mathfrak{s})$ to be the largest integer $n$ such that for any projective object $P$, there exist $n$ $\mathfrak{s}$-triangles
[…]
with $I^{k}$ being projective-injective for $10\leq k\leq n-$ 1. If such an $n$ does not exist, we let $\mathrm{dom.dim}(\mathcal{C},\mathbb{E},\mathfrak{s})=\infty$.
For an object $X$ in an extriangulated category $(\mathcal{C},\mathbb{E},\mathfrak{s})$ and $\mathcal{I}\subset\mathcal{C}$ a full additive subcategory consisting of injective objects, we say that $X$ admits an $\mathcal{I}$-coresolution of length $l$ if there exist $\mathfrak{s}$-triangles
[…]
where $I^{i}\in\mathcal{I}$ for $10\leq i\leq l-$ 1. We call $X^{i}$ the $i$-th cosyzygy of $X$ (with respect to $\mathcal{I}$).
Definition 2.2.
An extriangulated category $(\mathcal{C},\mathbb{E},\mathfrak{s})$ is $d$-minimal Auslander–Gorenstein if
  • (a)
    $\mathcal{C}$ has enough projectives;
  • (b)
    each projective has injective codimension $\leq d+1$, i.e. $\mathbb{E}^{d+2}(-,P)=0$ for each projective object $P$ in $\mathcal{C}$;
  • (c)
    it has dominant dimension $\geq d+1$.
An exact dg category $\mathcal{A}$ is $d$-minimal Auslander–Gorenstein if $H^{0}(\mathcal{A})$ is $d$-minimal Auslander–Gorenstein.
Remark 2.3.
A Frobenius extriangulated category is $d$-minimal Auslander–Gorenstein for each $d\geq 0$.
For a pretriangulated dg category $\mathcal{T}$ and full dg subcategories $\mathcal{X}$ and $\mathcal{Y}$, by abuse of notation, we write $\mathcal{X}\ast\mathcal{Y}$ as the full dg subcategory of $\mathcal{T}$ consisting of the objects $Z$ which admit triangles in the triangulated category $H^{0}(\mathcal{T})$
$X\rightarrow Z\rightarrow Y\rightarrow\Sigma X$
with $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$. For a triple $(\mathcal{K},\mathcal{P},\mathcal{I})$, where $\mathcal{K}$ is an additive dg category and $\mathcal{I}\subset\mathcal{P}$ are connective additive dg subcategories of $\mathcal{K}$, and for $n\geq 0$, we denote by $\mathcal{B}^{(n)}_{\mathcal{K},\mathcal{P},\mathcal{I}}$ the full dg subcategory of $\mathrm{pretr}(\mathcal{K})$ consisting of the objects in
$\mathcal{K}\ast\Sigma\mathcal{K}\ast\ldots\ast\Sigma^{n-1}\mathcal{K}\ast\Sigma^{n}\mathcal{K}\cap\mathrm{ker}\mathrm{Ext}^{\geq 1}(-,\mathcal{I}).$
A priori, the dg subcategory $\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(n)}$ is not stable under extensions in $\mathrm{pretr}(\mathcal{K})$. Below we will show that this is indeed the case when the triple $(\mathcal{K},\mathcal{P},\mathcal{I})$ is $d$-precluster tilting and $n\leq d+1$.
Definition 2.4.
Let $\mathcal{K}$ be an additive dg category and $\mathcal{I}\subseteq\mathcal{P}\subseteq\mathcal{K}$ full additive dg subcategories. The triple $(\mathcal{K},\mathcal{P},\mathcal{I})$ is $d$-precluster tilting if it has the following properties
  • (1)
    $\mathcal{P}$ is connective;
  • (2)
    $H^{0}(\mathcal{I})$ is covariantly finite in $H^{0}(\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d)})$;
  • (3)
    $\mathcal{K}$ is extension-closed in $\mathrm{tr}(\mathcal{K})$;
  • (4)
    $\mathrm{Hom}_{\mathrm{tr}(\mathcal{K})}(\mathcal{P},\Sigma^{\geq i}\mathcal{K})=0$ and $\mathrm{Hom}_{\mathrm{tr}(\mathcal{K})}(\mathcal{K},\Sigma^{\geq i}\mathcal{P})=0$ for any $i\geq 1$;
  • (5)
    for each object $K$ in $\mathcal{K}$, there is a triangle in $\mathrm{tr}(\mathcal{K})$
    $K^{\prime}\rightarrow P\rightarrow K\rightarrow\Sigma K^{\prime}$
    where $P\in\mathcal{P}$ and $K^{\prime}\in\mathcal{K}$.
Note that the dg category $\mathcal{K}$ is usually not connective and we could rewrite property $(5)$ in Definition 2.4 as follows: $\mathcal{K}\subset\mathcal{P}\ast\Sigma\mathcal{K}$ in $\mathrm{pretr}(\mathcal{K})$. We denote by $\mathcal{L}_{\mathcal{K},\mathcal{P},\mathcal{I}}$ (resp. $\mathcal{Q}_{\mathcal{K},\mathcal{P},\mathcal{I}}$, $\mathcal{J}_{\mathcal{K},\mathcal{P},\mathcal{I}}$) the full dg subcategory of $\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ on the objects in the closure under kernels of retractions of the objects in $H^{0}(\mathcal{K})$ (resp. in $H^{0}(\mathcal{P})$, $H^{0}(\mathcal{I})$). The dg category $\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ is a $d$-minimal Auslander–Gorenstein exact dg category as the following proposition shows.
Proposition 2.5.
Let $(\mathcal{K},\mathcal{P},\mathcal{I})$ be a $d$-precluster tilting triple of dg categories. Then $\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ is extension-closed in $\mathrm{pretr}(\mathcal{K})$. We endow it with the inherited exact dg structure. The following holds for $\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$.
  • (a)
    $\mathcal{Q}_{\mathcal{K},\mathcal{P},\mathcal{I}}$ is the full dg subcategory of projectives and $\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ has enough projectives;
  • (b)
    each projective has injective codimension $\leq d+1$;
  • (c)
    $\mathcal{J}_{\mathcal{K},\mathcal{P},\mathcal{I}}$ is the full dg subcategory of projective-injectives and $\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ has dominant dimension $\geq d+1$.
Proof.
Indeed, it is easy to verify that $\mathcal{P}\ast\Sigma\mathcal{P}\ast\ldots\ast\Sigma^{d}\mathcal{P}\ast\Sigma^{d+1}\mathcal{K}$ is extension-closed in $\mathrm{pretr}(\mathcal{K})$ by (3) and (4) in Definition 2.4: we have
$\Sigma^{d+1}\mathcal{K}\ast\Sigma^{i}\mathcal{P}\subseteq\Sigma^{i}\mathcal{P}\ast\Sigma^{d+1}\mathcal{K},\;\;i\leq d+1,$
and that it is equal to $\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$:
$\mathcal{K}\ast\Sigma\mathcal{K}\ast\ldots\ast\Sigma^{d}\mathcal{K}\ast\Sigma^{d+1}\mathcal{K}$
$\subset$ $(\mathcal{P}\ast\Sigma\mathcal{K})\ast\Sigma\mathcal{K}\ast\Sigma^{2}\mathcal{K}\ast\ldots\ast\Sigma^{d}\mathcal{K}\ast\Sigma^{d+1}\mathcal{K}$
$\subset$ $\mathcal{P}\ast\Sigma\mathcal{K}\ast\Sigma^{2}\mathcal{K}\ast\ldots\ast\Sigma^{d}\mathcal{K}\ast\Sigma^{d+1}\mathcal{K}$
$\subset$ $\ldots$
$\subset$ $\mathcal{P}\ast\Sigma\mathcal{P}\ast\ldots\ast\Sigma^{d}\mathcal{P}\ast\Sigma^{d+1}\mathcal{K}.$
Then it is clear that $\mathcal{Q}_{\mathcal{K},\mathcal{P},\mathcal{I}}$ is the full dg subcategory of projectives and $\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ has enough projectives. It is also clear that $\mathcal{J}_{\mathcal{K},\mathcal{P},\mathcal{I}}$ is the full dg subcategory of projective-injectives.
Item (b) follows from property (4) and the definition of $H^{0}(\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)})$.
Item (c) follows from property (2). Indeed, let $Q$ be an object in $\mathcal{Q}_{\mathcal{K},\mathcal{P},\mathcal{I}}$. Then we have objects $P$ and $S$ in $\mathcal{P}$ such that $P\xrightarrow{{}_{\sim}}Q\oplus S$. We have the following diagram in $\mathrm{tr}(\mathcal{K})$
[…]
where $f$ is a left $H^{0}(\mathcal{I})$-approximation. Now the claim follows from the Horseshoe Lemma applied to the conflation in the bottom row. ∎
For a connective exact dg category $\mathcal{A}$, we denote by $\overline{\mathcal{A}}$ the full dg subcategory of $\mathcal{D}^{b}_{dg}(\mathcal{A})$ consisting of the objects in the closure under kernels of retractions of $H^{0}(\mathcal{A})$ in $\mathcal{D}^{b}(\mathcal{A})$.
Definition 2.6 ([6, Definition 3.6]).
  1. 1)
    Two connective exact dg categories $\mathcal{A}$ and $\mathcal{B}$ are equivalent if there is a quasi-equivalence $\overline{\mathcal{A}}\xrightarrow{{}_{\sim}}\overline{\mathcal{B}}$.
  2. 2)
    Two $d$-precluster tilting triples $(\mathcal{K},\mathcal{P},\mathcal{I})$ and $(\mathcal{K}^{\prime},\mathcal{P}^{\prime},\mathcal{I}^{\prime})$ are equivalent if there is a quasi-equivalence $\mathcal{L}_{\mathcal{K},\mathcal{P},\mathcal{I}}\xrightarrow{{}_{\sim}}\mathcal{L}_{\mathcal{K}^{\prime},\mathcal{P}^{\prime},\mathcal{I}^{\prime}}$ which restricts to quasi-equivalences $\mathcal{Q}_{\mathcal{K},\mathcal{P},\mathcal{I}}\xrightarrow{{}_{\sim}}\mathcal{Q}_{\mathcal{K}^{\prime},\mathcal{P}^{\prime},\mathcal{I}^{\prime}}$ and $\mathcal{J}_{\mathcal{K},\mathcal{P},\mathcal{I}}\xrightarrow{{}_{\sim}}\mathcal{J}_{\mathcal{K}^{\prime},\mathcal{P}^{\prime},\mathcal{I}^{\prime}}$.
Remark 2.7.
For a $d$-precluster tilting triple $(\mathcal{K},\mathcal{P},\mathcal{I})$, we have that $\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ is stable under kernels of retractions in $\mathrm{pretr}(\mathcal{K})$. The proof is similar to that of [6, Proposition 3.10]. Via the quasi-equivalence $\mathrm{pretr}(\mathcal{K}_{\mathcal{A}})\rightarrow\mathcal{D}^{b}_{dg}(\mathcal{A})$, the canonical morphism $\mathcal{A}\rightarrow\tau_{\leq 0}\mathcal{B}_{\mathcal{K}_{A},\mathcal{P}_{\mathcal{A}},\mathcal{I}_{\mathcal{A}}}$ identifies $\tau_{\leq 0}\mathcal{B}_{\mathcal{K}_{A},\mathcal{P}_{\mathcal{A}},\mathcal{I}_{\mathcal{A}}}$ with $\overline{\mathcal{A}}$, cf. Lemma 2.9 (e).
Theorem 2.8.
There is a canonical bijection between the following
  • (1)
    equivalence classes of connective exact dg categories $\mathcal{A}$ which are $d$-minimal Auslander–Gorenstein;
  • (2)
    equivalence classes of $d$-precluster tilting triples $(\mathcal{K},\mathcal{P},\mathcal{I})$.
The bijection from $(1)$ to $(2)$ sends $\mathcal{A}$ to the triple $(\mathcal{K},\mathcal{P},\mathcal{I})$ formed by the full dg subcategory $\mathcal{P}$ on the projectives of $\mathcal{A}$, its full dg subcategory $\mathcal{I}$ of projective-injectives and the full dg subcategory $\mathcal{K}$ of $\mathcal{D}^{b}_{dg}(\mathcal{A})$ consisting of the objects $K$ which are $(d+1)$-st syzygies of objects in $\mathcal{A}$ (in particular we have $\mathrm{Ext}^{\geq 1}(K,\mathcal{P})=0$). The inverse bijection sends $(\mathcal{K},\mathcal{P},\mathcal{I})$ to the $\tau_{\leq 0}$-truncation of the dg subcategory $\mathcal{B}^{(d+1)}_{\mathcal{K},\mathcal{P},\mathcal{I}}$ of $\mathrm{pretr}(\mathcal{K})$.
We will prove Theorem 2.8 after Proposition 2.11.
Lemma 2.9.
Let $\mathcal{A}$ be a connective $d$-minimal Auslander–Gorenstein exact dg category. Let $\mathcal{P}$ (resp. $\mathcal{I}$) be the full dg subcategory on the projectives (resp. projective-injectives) of $\mathcal{A}$. Let $\mathcal{K}$ be the full dg subcategory of $\mathcal{D}^{b}_{dg}(\mathcal{A})$ consisting of the objects $K$, which are $(d+1)$-st syzygies of objects in $\mathcal{A}$. Then the following statements hold:
  • (a)
    The universal embedding $\mathcal{A}\rightarrow\mathcal{D}^{b}_{dg}(\mathcal{A})$ induces a quasi-fully faithful morphism $\mathcal{P}\hookrightarrow\mathcal{K}$. We view $\mathcal{P}$ as a full dg subcategory of $\mathcal{K}$ via this morphism.
  • (b)
    Each object $K$ in $\mathcal{K}$ admits triangles in $\mathcal{D}^{b}(\mathcal{A})$
[…]
    where $I^{i}\in\mathcal{I}$ and $L^{i+1}\in\mathcal{A}$ for $0\leq i\leq d$. In particular, by the Horseshoe Lemma, the category $H^{0}(\mathcal{K})$ is extension-closed in $H^{0}(\mathcal{D}^{b}_{dg}(\mathcal{A}))$.
  • (c)
    We have $\mathcal{K}\subseteq\mathcal{P}\ast\Sigma\mathcal{K}$ in $\mathrm{pretr}(\mathcal{K})$ and $\mathrm{Ext}^{\geq 1}(\mathcal{K},\mathcal{P})=0$ and $\mathrm{Ext}^{\geq 1}(\mathcal{P},\mathcal{K})=0$.
  • (d)
    $H^{0}(\mathcal{I})$ is covariantly finite in $H^{0}(\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d)})$.
  • (e)
    $\mathcal{A}$ is equivalent to $\tau_{\leq 0}\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ in the sense of Definition 2.6 1), and $\mathrm{pretr}(\mathcal{K})$ is quasi-equivalent to $\mathcal{D}^{b}_{dg}(\mathcal{A})$
In particular, the triple $(\mathcal{K},\mathcal{P},\mathcal{I})$ is $d$-precluster tilting.
Proof.
Statement (a) is straightforward.
Let $K$ be an object in $\mathcal{K}$. By definition, there exists triangles in $\mathcal{D}^{b}(\mathcal{A})$
[…]
where $P^{i}\in\mathcal{P}$ and $M^{i+1}\in\mathcal{A}$ for $0\leq i\leq d$. We prove by induction that the object $M^{i}$ has an $\mathcal{I}$-coresolution of length $d+1-i$ for $10\leq i\leq d+$ 1. Then statement (b) will follow from the case when $i=0$. For $i=d+1$, the statement is clearly true. We assume that the statement holds for $M^{i+1}$. Since $H^{0}(\mathcal{A})$ is $d$-minimal Auslander–Gorenstein, the object $P^{i}$ has an $\mathcal{I}$-coresolution of length $d+1$. So there is a triangle
$P^{i}\rightarrow I^{i}\rightarrow U^{i}\rightarrow\Sigma P^{i}$
where $U^{i}$ has $\mathcal{I}$-coresolution of length $d$ and we have the following diagram in $\mathcal{D}^{b}(\mathcal{A})$
[…]
By the Horseshoe Lemma, the object $V^{i}$ has $\mathcal{I}$-coresolution of length $d-i$. It follows that the object $M^{i}$ has $\mathcal{I}$-coresolution of length $d+1-i$ and this finishes the proof of the induction step.
Let us show statement (c). We have a quasi-fully faithful morphism $\mathrm{pretr}(\mathcal{K})\hookrightarrow\mathcal{D}^{b}_{dg}(\mathcal{A})$. Therefore it is enough to show that $\mathcal{K}\subseteq\mathcal{P}\ast\Sigma\mathcal{K}$ in $\mathcal{D}^{b}_{dg}(\mathcal{A})$. Let $K$ be an object in $\mathcal{K}$ with triangles (2.9.12.9.3) where $P^{i}\in\mathcal{P}$ and $M^{i+1}\in\mathcal{A}$ for $0\leq i\leq d$. Since $H^{0}(\mathcal{A})$ has enough projectives, we have a triangle in $\mathcal{D}^{b}(\mathcal{A})$
[…]
Then the triangle (2.9.4) and triangles (2.9.12.9.2) show that $K^{\prime}\in\mathcal{K}$ and hence $\mathcal{K}\subseteq\mathcal{P}\ast\Sigma\mathcal{K}$. From the triangles (2.9.12.9.3), we have $\mathrm{Ext}^{i}(M^{j},P)\xrightarrow{{}_{\sim}}\mathrm{Ext}^{i+1}(M^{j+1},P)$ for $i\geq 1$ and $P\in\mathcal{P}$ and $j\geq 0$. It follows that $\mathrm{Ext}^{i}(K,P)\xrightarrow{{}_{\sim}}\mathrm{Ext}^{i+d+1}(M^{d+1},P)$ for $i\geq 1$ and $P\in\mathcal{P}$. Since the injective codimensions of projectives are $\leq d+1$, it follows that $\mathrm{Ext}^{\geq 1}(K,P)=0$ for $P\in\mathcal{P}$. Since $\mathcal{P}$ consists of projective objects in $\mathcal{A}$, we also have $\mathrm{Ext}^{\geq 1}(P,K)=0$ for $P\in\mathcal{P}$.
Let us show statement (e). We have the following diagram
[…]
Indeed, we need to show
  • (1)
    the canonical morphism $\mathrm{pretr}(\mathcal{K})\hookrightarrow\mathcal{D}^{b}_{dg}(\mathcal{A})$ is quasi-dense, and
  • (2)
    $\mathcal{A}$ is contained in $\mathcal{B}=\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$, and
  • (3)
    under the quasi-fully faithful morphism $\mathcal{A}\rightarrow\tau_{\leq 0}\mathcal{B}$, each object in $\mathcal{B}$ is the kernel of a retraction in $\mathcal{A}$.
For each object $X$ in $\mathcal{A}$, we have conflations in $H^{0}(\mathcal{A})$
[…]
where $P_{i}\in H^{0}(\mathcal{P})$ for each $0\leq i\leq d$. It follows that $K_{d+1}\in\mathcal{K}$. This shows item (2). Since $\mathcal{P}$ is contained in $\mathcal{K}$, item (1) also follows.
Let us show item (3). For an object $X$ in $\mathcal{B}$, it admits conflations (2.9.5) where $P_{i}$ is projective for each $0\leq i\leq d$ and $K_{d+1}\in\mathcal{K}$. Let us show that $X$ is in the closure of kernels of retractions in $\mathcal{A}$. We prove by induction on $i$ that $X_{i}$ admits an $\mathcal{I}$-coresolution of length $i$ with the $i$-th cosygyzy an object in $\mathcal{A}$ for each $10\leq i\leq d+$ 1. It obviously holds for $X_{d+1}=K_{d+1}\in H^{0}(\mathcal{A})$: By item (b) the object $K_{d+1}$ admits an $\mathcal{I}$-coresolution of length $d+1$ with the $(d+1)$-cosyzygy an object in $\mathcal{A}$.
Let us assume that it holds for $i+1\geq 1$. So we have a conflation
$X_{i+1}\rightarrow I_{i+1}\rightarrow U_{i+1}$
where $I_{i+1}\in\mathcal{I}$ and $U_{i+1}\in H^{0}(\mathcal{A})$ admits an $\mathcal{I}$-coresolution of length $i$ with the $i$-th cosygyzy an object in $\mathcal{A}$. We have the following diagram
[…]
From the second column, it follows by the Horseshoe Lemma that $V_{i+1}\xrightarrow{{}_{\sim}}X_{i}\oplus I_{i+1}\in H^{0}(\mathcal{A})$ admits an $\mathcal{I}$-coresolution of length $i$ with the $i$-th cosygyzy an object in $\mathcal{A}$. A similar argument to that of Proposition 2.5 (c) shows that the same is true for $X_{i}$. This finishes the proof of the induction step. We apply the above argument to the case $i=0$ and it follows that $X\oplus I_{1}\xrightarrow{{}_{\sim}}V_{1}\in\mathcal{A}$ and hence $X$ is in the closure under kernels of retractions in $H^{0}(\mathcal{A})$. This shows item (3) and we are done.
Let us show statement (d). Let $N$ be an object in $\mathcal{B}^{(d)}_{\mathcal{K},\mathcal{P},\mathcal{I}}$. By definition, it admits triangles in $\mathrm{tr}(\mathcal{K})$
[…]
where $K^{i}\in\mathcal{K}$ for $0\leq i\leq d$ and $N^{j}\in\mathcal{B}^{(d)}_{\mathcal{K},\mathcal{P},\mathcal{I}}$ for $1\leq j\leq d$. By item (b), the object $K^{0}\in\mathcal{K}$ admits an $\mathcal{I}$-coresolution of length $d+1$. So we have the following diagram in $\mathrm{tr}(\mathcal{K})$
[…]
where $V$ admits an $\mathcal{I}$-coresolution of length $d$. By the Horseshoe Lemma, the object $U\xrightarrow{{}_{\sim}}N^{1}\oplus I$ also admits an $\mathcal{I}$-coresolution of length $d$. So we replace $N^{1}$ by $N^{1}\oplus I$ and we repeat the above argument. We see that $N=N^{d}$ admits an inflation into an object in $\mathcal{I}$. Hence $H^{0}(\mathcal{I})$ is covariantly finite in $H^{0}(\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d)})$. ∎
For a $d$-precluster tilting triple $(\mathcal{K},\mathcal{P},\mathcal{I})$, put $\mathcal{B}=\mathcal{B}^{(d+1)}_{\mathcal{K},\mathcal{P},\mathcal{I}}$ and $\mathcal{B}^{\prime}=\tau_{\leq 0}\mathcal{B}$.
Lemma 2.10.
Let $(\mathcal{K},\mathcal{P},\mathcal{I})$ be a $d$-precluster tilting triple. Then we have a canonical quasi-equvalence $\mathcal{D}^{b}_{dg}(\mathcal{B}^{\prime})\xrightarrow{{}_{\sim}}\mathrm{pretr}(\mathcal{K})$.
Proof.
By Proposition 2.5, the dg category $\mathcal{B}$ is stable under extensions in $\mathrm{pretr}(\mathcal{K})$ and inherits a canonical exact structure. Let $(H^{0}(\mathcal{B}),\mathbb{E},\mathfrak{s})$ be the canonical extriangulated structure.
We have the following diagram in $\mathrm{Hqe}$
[…]
Let $X$ be an object in $\mathcal{B}^{\prime}$. Then we have a conflation in $H^{0}(\mathcal{B}^{\prime})$
(2.10.1) $X_{1}\rightarrow P_{0}\rightarrow X$
where $P_{0}\in H^{0}(\mathcal{P})$. Let $Y$ be any object in $\mathcal{B}^{\prime}$. Then we have
$\mathrm{Hom}_{\mathcal{D}^{b}(\mathcal{B}^{\prime})}(X,\Sigma Y)\xrightarrow{{}_{\sim}}\mathrm{Hom}_{\mathrm{tr}(\mathcal{K})}(X,\Sigma Y)$
since both are isomorphic to $\mathbb{E}(X,Y)$. The conflation (2.10.1) gives rise to triangles in $\mathrm{tr}(\mathcal{K})$ and in $\mathcal{D}^{b}(\mathcal{B}^{\prime})$. We apply the functor $\mathrm{Hom}_{\mathcal{D}^{b}(\mathcal{B}^{\prime})}(-,Y)$ to the triangle in $\mathcal{D}^{b}(\mathcal{B}^{\prime})$ corresponding to the conflation (2.10.1) and we obtain a long exact sequence
$\mathrm{Hom}(X,Y)$ $\rightarrow\mathrm{Hom}(P_{0},Y)\rightarrow\mathrm{Hom}(X_{1},Y)\rightarrow\mathrm{Ext}^{1}(X,Y)$
$\rightarrow\mathrm{Ext}^{1}(P_{0},Y)=0\rightarrow\mathrm{Ext}^{1}(X_{1},Y)\rightarrow\mathrm{Ext}^{2}(X,Y)\rightarrow\mathrm{Ext}^{2}(P_{0},Y)=0\rightarrow\cdots$
We also have a corresponding long exact sequence in $\mathrm{pretr}(\mathcal{K})$. Note that $\mathrm{Ext}^{\geq 1}_{\mathcal{D}^{b}(\mathcal{B}^{\prime})}(P_{0},Y)$ vanishes since $P_{0}$ is projective in $\mathcal{B}^{\prime}$ and $\mathrm{Ext}^{\geq 1}_{\mathrm{tr}(\mathcal{K})}(P_{0},Y)$ vanishes by the definition of $d$-precluster tilting triple. It follows that
$\mathrm{Hom}_{\mathcal{D}^{b}(\mathcal{B}^{\prime})}(X,\Sigma^{2}Y)\xrightarrow{{}_{\sim}}\mathrm{Hom}_{\mathrm{tr}(\mathcal{K})}(X,\Sigma^{2}Y)$
and similarly for higher extension groups. It follows that $\mathrm{pretr}(\mathcal{K})$ is quasi-equivalent to $\mathcal{D}^{b}_{dg}(\mathcal{B}^{\prime})$. ∎
Proposition 2.11.
Let $(\mathcal{K},\mathcal{P},\mathcal{I})$ be a $d$-precluster tilting triple. We identify $\mathrm{pretr}(\mathcal{K})$ with $\mathcal{D}^{b}_{dg}(\mathcal{B}^{\prime})$ by Lemma 2.10. Then the closure of $\mathcal{K}$ under kernels of retractions in $\mathcal{D}^{b}_{dg}(\mathcal{B}^{\prime})$ is the full dg subcategory of $\mathcal{D}^{b}_{dg}(\mathcal{B}^{\prime})$ consisting of the objects $K$, which are (d+1)-th syzygies of the objects in $\mathcal{B}^{\prime}$.
Proof.
By Proposition 2.5, the closure $\mathcal{Q}_{\mathcal{K},\mathcal{P},\mathcal{I}}$ (resp. $\mathcal{J}_{\mathcal{K},\mathcal{P},\mathcal{I}}$) of the dg category $\mathcal{P}$ (resp. $\mathcal{I}$) under kernels of retractions is the full dg subcategory of $\mathcal{B}^{\prime}$ on the projectives (resp. projective-injectives).
Let $K$ be an object in $\mathcal{D}^{b}(\mathcal{B}^{\prime})$ which admits triangles (2.9.12.9.3) where $P^{i}\in\mathcal{Q}_{\mathcal{K},\mathcal{P},\mathcal{I}}$ and $M^{i+1}\in\mathcal{B}^{\prime}$ for $0\leq i\leq d$. We may assume that $\mathcal{P}^{i}$ lies in $\mathcal{P}$ for $0\leq i\leq d$. By definition of $\mathcal{B}^{\prime}$, the object $M^{d+1}$ admits triangles in $\mathcal{D}^{b}(\mathcal{B}^{\prime})$
[…]
where $Q^{i}\in\mathcal{P}$ and $N^{i}\in\mathcal{B}^{\prime}$ for $0\leq i\leq d$, and $K^{\prime}\in\mathcal{K}$. Since objects in $\mathcal{P}$ are projective in $H^{0}(\mathcal{B}^{\prime})$, we have the following diagram
[…]
where the sequence
[…]
is a conflation and is thus split. We repeat the argument and we see that we have an isomorphism $K^{\prime}\oplus P\xrightarrow{{}_{\sim}}K\oplus Q$ for some objects $P$ and $Q$ in $\mathcal{P}$. Hence $K$ is an object in the closure of $\mathcal{K}$ under kernels of retractions.
Conversely, let $M$ be an object in $\mathcal{D}^{b}_{dg}(\mathcal{B}^{\prime})$ such that $M\oplus K^{\prime}\xrightarrow{{}_{\sim}}K$ for some objects $K$ and $K^{\prime}$ in $\mathcal{K}$. By property (2) of Definition 2.4, the subcategory $H^{0}(\mathcal{I})$ is covariantly finite in $H^{0}(\mathcal{B}^{(d)}_{\mathcal{K},\mathcal{P},\mathcal{I}})$. Hence the objects in $\mathcal{K}$ are (d+1)-th syzygies of the objects in $\mathcal{B}^{\prime}$. We have the following diagram in $\mathcal{D}^{b}(\mathcal{B}^{\prime})$
[…]
where $V$ admits an $\mathcal{I}$-coresolution of length $d$. By the Horseshoe Lemma, the object $U$ also admits an $\mathcal{I}$-coresolution of length $d$ and hence $M$ is a (d+1)-th syzygy of the objects in $\mathcal{B}^{\prime}$.
Proof of Theorem 2.8.
Let $\mathcal{A}$ be a connective $d$-minimal Auslander–Gorenstein exact dg category. Let $(\mathcal{K}_{\mathcal{A}},\mathcal{P}_{\mathcal{A}},\mathcal{I}_{\mathcal{A}})$ be the triple formed by the full dg subcategory $\mathcal{P}_{\mathcal{A}}$ on the projectives of $\mathcal{A}$ and its full dg subcategory $\mathcal{I}_{\mathcal{A}}$ of projective-injectives and the full dg subcategory $\mathcal{K}_{\mathcal{A}}$ of $\mathcal{D}^{b}_{dg}(\mathcal{A})$ consisting of the objects $K$ which are $(d+1)$-st syzygies of objects in $\mathcal{A}$. By Lemma 2.9, the triple $(\mathcal{K}_{\mathcal{A}},\mathcal{P}_{\mathcal{A}},\mathcal{I}_{\mathcal{A}})$ is $d$-precluster tilting. Therefore the map sending $\mathcal{A}$ to $(\mathcal{K}_{\mathcal{A}},\mathcal{P}_{\mathcal{A}},\mathcal{I}_{\mathcal{A}})$ is well-defined.
Let $(\mathcal{K},\mathcal{P},\mathcal{I})$ be a $d$-precluster tilting triple. By Proposition 2.5, the connective exact dg category $\tau_{\leq 0}\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ is $d$-minimal Auslander–Gorenstein. Therefore the map sending $(\mathcal{K},\mathcal{P},\mathcal{I})$ to $\tau_{\leq 0}\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ is well-defined.
By Lemma 2.9 (e), for a connective $d$-minimal Auslander–Gorenstein exact dg category $\mathcal{A}$, we have that $\mathcal{A}$ and $\tau_{\leq 0}\mathcal{B}_{\mathcal{K}_{\mathcal{A}},\mathcal{P}_{\mathcal{A}},\mathcal{I}_{\mathcal{A}}}^{(d+1)}$ are equivalent in the sense of Definition 2.6.
Let $(\mathcal{K},\mathcal{P},\mathcal{I})$ be a $d$-precluster tilting triple. Put $\mathcal{B}^{\prime}=\tau_{\leq 0}\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$. By Proposition 2.5 and Proposition 2.11, we have a quasi-equivalence $\mathrm{pretr}(\mathcal{K})\xrightarrow{{}_{\sim}}\mathcal{D}^{b}_{dg}(\mathcal{B}^{\prime})$ which shows that the $d$-precluster tilting triples $(\mathcal{K},\mathcal{P},\mathcal{I})$ and $(\mathcal{K}_{\mathcal{B}^{\prime}},\mathcal{P}_{\mathcal{B}^{\prime}},\mathcal{I}_{\mathcal{B}^{\prime}})$ are equivalent. ∎
Example 2.12.
Let $k$ be a field. Let $\mathcal{K}$ be an extension-closed subcategory of a pretriangulated dg category $\mathcal{T}_{dg}$. Then $\mathcal{K}$ inherits a canonical exact structure which we assume to be Frobenius. Let $\mathcal{P}$ be the full dg subcategory of $\mathcal{K}$ consisting of the projectives and $\mathcal{I}\subset\mathcal{P}$ an additive full dg subcategory which is covariantly finite in $\mathcal{T}=H^{0}(\mathcal{T}_{dg})$. We assume that $\mathcal{P}$ is a connective dg category. Then the triple $(\mathcal{K},\mathcal{P},\mathcal{I})$ is $d$-precluster tilting for any $d\geq 0$. Indeed, for each object $K$ in $\mathcal{K}$, we have a triangle in $\mathcal{T}$
$K^{\prime}\rightarrow P\rightarrow K\rightarrow\Sigma K^{\prime}$
where $P\in\mathcal{P}$ and $K^{\prime}\in\mathcal{K}$, since $\mathcal{K}$ is Frobenius. Then we have
$\mathrm{Hom}_{\mathcal{T}}(Q,\Sigma^{i}K)\xrightarrow{{}_{\sim}}\mathrm{Hom}_{\mathcal{T}}(Q,\Sigma^{i+1}K^{\prime})$
for $i\geq 1$ and $Q\in\mathcal{P}$, by the assumption that $\mathcal{P}$ is connective. Since $\mathcal{K}$ is Frobenius, we have that $\mathrm{Hom}_{\mathcal{T}}(Q,\Sigma K)=0$ for $Q\in\mathcal{P}$ and $K\in\mathcal{K}$. Therefore we have that $\mathrm{Hom}_{\mathcal{T}}(Q,\Sigma^{i}K)=0$ for $i\geq 1$ and $Q\in\mathcal{P}$ and $K\in\mathcal{K}$. Similarly we have that $\mathrm{Hom}_{\mathcal{T}}(K,\Sigma^{i}Q)=0$ for $i\geq 1$ and $Q\in\mathcal{P}$ and $K\in\mathcal{K}$.
We have the following two classes of examples. The first class of examples is taken from [18]. Let $(Q,F,W)$ be a Jacobi-finite ice quiver with potential. Denote by $\Gamma_{rel}$ the relative Ginzburg dg algebra $\Gamma_{rel}(Q,F,W)$, which is a connective dg $k$-algebra. Let $e=\sum_{i\in F}e_{i}$ be the idempotent associated with all frozen vertices. Let $\mathrm{pvd}_{e}(\Gamma_{rel})$ be the full subcategory of $\mathrm{pvd}(\Gamma_{rel})$ of the dg $\Gamma_{rel}$-modules whose restriction to frozen vertices is acyclic. Then the relative cluster category $\mathcal{C}=\mathcal{C}(Q,F,W)$ associated to $(Q,F,W)$ is defined as the Verdier quotient of triangulated categories
$\mathrm{per}(\Gamma_{rel})/\mathrm{pvd}_{e}(\Gamma_{rel}).$
Let $\pi^{rel}:\mathrm{per}(\Gamma_{rel})\rightarrow\mathcal{C}$ be the quotient functor. Since $\mathrm{Hom}_{\mathrm{per}\Gamma_{rel}}(\mathcal{P},\mathrm{pvd}_{e}(\Gamma_{rel}))=0$, the functor $\pi^{rel}$ restricted to the thick category generated by $\mathcal{P}$ is fully faithful. By [18, Theorem 5.42], the Higgs category $\mathcal{H}$ is equal to the intersection ${{}^{\perp_{>0}}\mathcal{P}\cap\mathcal{P}^{\perp_{>0}}}$ in $\mathcal{C}$. In particular, it is extension-closed in $\mathcal{C}$. Let $\mathcal{I}\subseteq\mathcal{P}$ be any full additive dg subcategory. We take canonical dg enhancements of the above categories, where we add the subscripts $dg$ correspondingly. Then $(\mathcal{H}_{dg},\mathcal{P}_{dg},\mathcal{I}_{dg})$ is $d$-precluster tilting for $d\geq 0$.
The second class is taken from [16]. Let $A$ a dg $k$-algebra. Recall that
  • $A$ is proper if $\sum_{i\in\mathbb{Z}}\dim H^{i}(A)<\infty$;
  • $A$ is Gorenstein if the perfect derived category $\mathrm{per}(A)$ coincides with the thick subcategory of $\mathcal{D}(A)$ generated by $DA$, where $D=\mathrm{Hom}(-,k)$ is the $k$-dual.
Let $A$ be a connective proper Gorenstein dg algebra. We denote by $\mathrm{pvd}(A)$ the full subcategory of $\mathcal{D}(A)$ on the perfectly valued dg $A$-modules, i.e. dg $A$-modules whose total cohomology is finite-dimensional. A dg $A$-module $M\in\mathrm{pvd}(A)$ is Cohen–Macaulay [16, Definition 2.1] if it is connective and if $\mathrm{Hom}_{\mathcal{D}(A)}(M,\Sigma^{i}A)=0$ for $i>0$. We denote by $\mathrm{CM}(A)$ the full subcategory of $\mathrm{pvd}(A)$ on the Cohen–Macaulay dg $A$-modules. Let $I$ be a direct summand of $A$. Let $\mathcal{K}=\mathrm{CM}_{dg}(A)$ be the canonical dg enhancement of $\mathrm{CM}(A)$ and $\mathcal{P}=\mathrm{add}(A)$ and $\mathcal{I}=\mathrm{add}(I)$ the corresponding full dg subcategories. Then the triple $(\mathcal{K},\mathcal{P},\mathcal{I})$ is $d$-precluster tilting for each $d\geq 0$. Let $\mathrm{pvd}^{\leq 0}(A)$ be the full subcategory of $\mathrm{pvd}(A)$ consisting of the dg $A$-modules whose cohomology is concentrated in degrees $\leq i$. The $d$-minimal Auslander–Gorenstein exact dg category $\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ is a full dg subcategory of $\mathrm{pvd}_{dg}^{\leq 0}(A)$. When $I=0$, we have that $\mathrm{pvd}_{dg}^{\leq 0}(A)$ is the union of $\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ where $d$ runs through non-negative integers, cf. [16, Proof of Theorem 3.7].
3. Connection to exact categories
In this section we give necessary and sufficient conditions on a $d$-precluster tilting triple $(\mathcal{K},\mathcal{P},\mathcal{I})$ for the connective $d$-minimal Aulander–Gorenstein exact dg category $\tau_{\leq 0}\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ to be concentrated in degree zero (resp. to be an abelian category). In the abelian case, we will see that $\mathcal{K}$ and $\mathcal{I}$ are determined by $\mathcal{P}$ and that there is a fully faithful functor $\mathcal{P}\rightarrow(\mathcal{I}\mbox{-}\mathrm{mod})^{op}$. Assume $k$ is a commutative artinian ring. When $\tau_{\leq 0}\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ is the module category over an Artin $k$-algebra $\Gamma$, we may identify $\mathcal{I}$ with the full subcategory of injective modules over some Artin $k$-algebra $\Lambda$ and identify $\mathcal{P}$ with an additive full subcategory of $\mathrm{mod}\Lambda$. Suppose that $\mathcal{P}=\mathrm{add}M$ for some $\Lambda$-module $M$. We show that the thus defined triple $(\mathcal{K},\mathcal{P},\mathcal{I})$ is $d$-precluster tilting if and only if $M$ is a $d$-precluster tilting $\Lambda$-module. This justifies our terminology to some extent.
3.1. Exact case
Proposition 3.2.
Let $(\mathcal{K},\mathcal{P},\mathcal{I})$ be a $d$-precluster tilting triple. The connective exact dg category $\tau_{\leq 0}\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ is concentrated in degree zero if and only if the following conditions hold
  • (a)
    The dg category $\mathcal{K}$ is concentrated in non-negative degrees (in particular, the dg category $\mathcal{P}$ is concentrated in degree zero).
  • (b)
    Consider a complex in $H^{0}(\mathcal{K})$
[…]
    where $K_{d+1}\in H^{0}(\mathcal{K})$ and $P_{i}\in H^{0}(\mathcal{P})$ for $0\leq i\leq d$. If the induced complex of Hom spaces in $H^{0}(\mathcal{K})$
[…]
    is exact for each $I\in\mathcal{I}$, then the corresponding complex of Hom spaces in $H^{0}(\mathcal{K})$
[…]
    is exact for each $P\in\mathcal{P}$.
Proof.
Let $(\mathcal{K},\mathcal{P},\mathcal{I})$ be a $d$-precluster tilting triple satisfying conditions (a) and (b). Put $\mathcal{B}=\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ and $\mathcal{B}^{\prime}=\tau_{\leq 0}\mathcal{B}$. Let $X$ be an object in $\mathcal{B}$. We have conflations in $H^{0}(\mathcal{B}^{\prime})$
[…]
where $K_{d+1}\in\mathcal{K}$ and $P_{i}\in\mathcal{P}$ for $0\leq i\leq d$. By condition (a), we have $\mathrm{Hom}_{\mathrm{tr}(\mathcal{K})}(\mathcal{K},\Sigma^{-i}\mathcal{P})=0$ for $i\geq 1$ and therefore $\mathrm{Hom}_{\mathrm{tr}(\mathcal{K})}(X_{j},\Sigma^{-i}\mathcal{P})=0$ for $i\geq 1$ and $0\leq j\leq d$. It follows that the complex (3.2.2) is exact for each $I\in\mathcal{I}$ and hence by condition (b), the complex (3.2.3) is exact for $P\in\mathcal{P}$. Since objects in $\mathcal{P}$ are projective in $H^{0}(\mathcal{B}^{\prime})$, the maps $P_{i}\rightarrow X_{i}$ induce surjections $\mathrm{Hom}(P,P_{i})\rightarrow\mathrm{Hom}(P,X_{i})$ for $P\in\mathcal{P}$ and $0\leq i\leq d$. Therefore the maps $\mathrm{Hom}(P,X_{i+1})\rightarrow\mathrm{Hom}(P,P_{i})$ are injections for $0\leq i\leq d$ and $P\in\mathcal{P}$. It follows that we have $\mathrm{Hom}(P,\Sigma^{-i}X_{j})=0$ for $P\in\mathcal{P}$ and $i\geq 1$ and $10\leq j\leq d+$ 1. Let $X^{\prime}$ be another object in $\mathcal{B}^{\prime}$ with conflations given by adding a prime symbol to each object in the conflations (3.2.4). Then we have
$\mathrm{Hom}(X^{\prime},\Sigma^{-1}X)\xrightarrow{{}_{\sim}}\mathrm{Hom}(X_{1}^{\prime},\Sigma^{-2}X)\xrightarrow{{}_{\sim}}\ldots\xrightarrow{{}_{\sim}}\mathrm{Hom}(K_{d+1},\Sigma^{-d-2}X)$
and
$\mathrm{Hom}(K,\Sigma^{-d-2}X)\xrightarrow{{}_{\sim}}\mathrm{Hom}(K,\Sigma^{-d-1}X_{1})\xrightarrow{{}_{\sim}}\ldots\xrightarrow{{}_{\sim}}\mathrm{Hom}(K,\Sigma^{-1}K_{d+1})=0$
for $K\in\mathcal{K}$. Similarly we have $\mathrm{Hom}(X^{\prime},\Sigma^{-i}X)=0$ for $i\geq 1$. Therefore the connective exact dg category $\mathcal{B}^{\prime}$ is concentrated in degree zero.
Conversely, we suppose that $\mathcal{B}^{\prime}$ is concentrated in degree zero. Since $\tau_{\leq 0}\mathcal{K}$ is a full dg subcategory of $\mathcal{B}^{\prime}$, we have that $\tau_{\leq 0}\mathcal{K}$ is concentrated in degree zero and hence $\mathcal{K}$ is concentrated in non-negative degrees. This shows condition (a).
Let us show condition (b). Since $\mathcal{B}^{\prime}$ is concentrated in degree zero, a complex (3.2.1) in $H^{0}(\mathcal{K})$ such that the corresponding complex (3.2.2) is exact for each $I\in\mathcal{I}$, where $K_{d+1}\in H^{0}(\mathcal{K})$ and $P_{i}\in H^{0}(\mathcal{P})$ for $0\leq i\leq d$, is equivalent to a sequence of conflations (3.2.4) in $H^{0}(\mathcal{B}^{\prime})$. Let $P$ be an object in $\mathcal{P}$. If we apply the functor $\mathrm{Hom}(P,-)$ to the sequences in (3.2.4), they become short exact sequences of Hom spaces in $H^{0}(\mathcal{K})$. Now the exactness of the complex (3.2.3) follows by pasting the above short exact sequences. ∎
Lemma 3.3.
Let $(\mathcal{K},\mathcal{P},\mathcal{I})$ be a $d$-precluster tilting triple which satisfies the conditions (a) and (b) in Proposition 3.2. Then the map sending $X\in\mathcal{B}^{\prime}=\tau_{\leq 0}\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ to $\mathrm{Hom}_{\mathcal{B}^{\prime}}(-,X)|_{\mathcal{P}}$ extends to a fully faithful exact functor $F:\mathcal{B}^{\prime}\hookrightarrow\mathrm{Mod}\,\mathcal{P}$. It induces a bijection
$\mathrm{Ext}^{1}_{\mathcal{B}^{\prime}}(X,X^{\prime})\xrightarrow{{}_{\sim}}\mathrm{Ext}^{1}_{\mathrm{Mod}\,\mathcal{P}}(FX,FX^{\prime})$
for $X$ and $X^{\prime}$ in $\mathcal{B}^{\prime}$.
Proof.
By Proposition 3.2, the exact dg category $\mathcal{B}^{\prime}$ is concentrated in degree zero and hence is an exact category. Since the objects in $\mathcal{P}$ are projective in $\mathcal{B}^{\prime}$, the functor $F$ is an exact functor.
Let $X$ be an object in $\mathcal{B}^{\prime}$ with conflations (3.2.4). Recall that for an object $P\in\mathcal{P}$, we write $P^{\wedge}$ for the representable right $\mathcal{P}$-module $\mathrm{Hom}_{\mathcal{P}}(-,P)$. The right $\mathcal{P}$-module $\mathrm{Hom}_{\mathcal{B}^{\prime}}(-,X)|_{\mathcal{P}}$ is the cokernel of the morphism $P_{1}^{\wedge}\rightarrow P_{0}^{\wedge}$. Let $X^{\prime}$ be another object in $\mathcal{B}^{\prime}$. Then we have the following diagram
[…]
where the second row is a sequence of Hom spaces in $\mathrm{Mod}\,\mathcal{P}$. Hence we have $\mathrm{Hom}_{\mathcal{B}^{\prime}}(X,X^{\prime})\xrightarrow{{}_{\sim}}\mathrm{Hom}_{\mathrm{Mod}\,\mathcal{P}}(FX,FX^{\prime})$ and the functor $F$ is fully faithful. Since objects in $\mathcal{P}$ (resp. $F(\mathcal{P})$) are projective in $\mathcal{B}^{\prime}$ (resp. $\mathrm{Mod}\,\mathcal{P}$), we have a natural bijection
$\mathrm{Ext}^{1}_{\mathcal{B}^{\prime}}(X,X^{\prime})\xrightarrow{{}_{\sim}}\mathrm{Ext}^{1}_{\mathrm{Mod}\,\mathcal{P}}(FX,FX^{\prime}).$
3.4. Abelian case
Proposition 3.5.
Let $(\mathcal{K},\mathcal{P},\mathcal{I})$ be a $d$-precluster tilting triple. The connective exact dg category $\mathcal{B}^{\prime}=\tau_{\leq 0}\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ is abelian if and only if the following conditions hold
  • (a)
    The dg category $\mathcal{K}$ is concentrated in non-negative degrees (in particular, the dg category $\mathcal{P}$ is concentrated in degree zero) and the category $\mathcal{P}$ admits weak kernels.
  • (b)
    Consider a complex in $H^{0}(\mathcal{K})$
[…]
    where $K_{d+1}\in H^{0}(\mathcal{K})$ and $P_{i}\in H^{0}(\mathcal{P})$ for $0\leq i\leq d$. The induced complex of Hom spaces in $H^{0}(\mathcal{K})$
[…]
    is exact for each $I\in\mathcal{I}$, if and only if the corresponding complex of Hom spaces in $H^{0}(\mathcal{K})$
[…]
    is exact for each $P\in\mathcal{P}$.
  • (c)
    Each morphism $P_{1}\rightarrow P_{0}$ in $H^{0}(\mathcal{P})$ can be completed to a complex of the form (3.5.1) where $K_{d+1}\in H^{0}(\mathcal{K})$ and $P_{i}\in H^{0}(\mathcal{P})$ for $0\leq i\leq d$ such that the corresponding complex (3.5.3) is exact for each $P\in\mathcal{P}$. In other words, the functor $F:\mathcal{B}^{\prime}\hookrightarrow\mathrm{mod}\mathcal{P}$ defined in Lemma 3.3 is dense.
Proof.
Let $(\mathcal{K},\mathcal{P},\mathcal{I})$ be a $d$-precluster tilting triple which satisfies conditions (a), (b) and (c) above. Let us show that $\mathcal{B}^{\prime}$ is abelian. By Lemma 3.3, we have an exact embedding $F:\mathcal{B}^{\prime}\hookrightarrow\mathrm{mod}\mathcal{P}$ which induces a bijection between the $\mathrm{Ext}^{1}(-,-)$. By condition (c), the functor $F$ is dense and hence $F$ is an equivalence of categories. It follows that $\mathcal{B}^{\prime}\xrightarrow{{}_{\sim}}\mathrm{mod}\mathcal{P}$ is an abelian category.
Conversely we assume that $\mathcal{B}^{\prime}$ (with the induced exact structure from $\mathrm{tr}(\mathcal{K})$) is an abelian category. Since $\mathcal{B}^{\prime}$ has enough projectives, it follows that $\mathcal{P}$ has weak kernels. This proves condition (a). Let us prove the “if” part of condition (b). For each object $Y$ in $\mathcal{B}^{\prime}$, there is a conflation
[…]
where $Q_{0}\in\mathcal{P}$. Then the morphism $Q_{0}\rightarrow Y$ is an epimorphism. Since the morphism $(P,K_{d+1})\rightarrowtail(P,P_{d})$ is an injection for each $P\in\mathcal{P}$, the morphism $K_{d+1}\rightarrow P_{d}$ is a monomorphism in $\mathcal{B}^{\prime}$. Therefore we have a triangle in $\mathrm{tr}(\mathcal{K})$
[…]
where $X_{d}\in\mathcal{B}^{\prime}$. Since $\mathrm{Hom}_{\mathrm{tr}(\mathcal{K})}(\Sigma K_{d+1},P_{d-1})=0$, we have a unique map $X_{d}\rightarrow P_{d-1}$ such that the following diagram commutes
[…]
A similar argument shows that the morphism $X_{d}\rightarrow P_{d-1}$ is a monomorphism in $\mathcal{B}^{\prime}$. We repeat the above procedures and we see that we have a sequence of conflations (3.2.4) in $H^{0}(\mathcal{B}^{\prime})$. By definition, we see that the complex 3.5.2 is exact for each $I\in\mathcal{I}$.
By Lemma 3.3, we have an exact embedding $F:\mathcal{B}^{\prime}\hookrightarrow\mathrm{mod}\mathcal{P}$ and the representable modules $P^{\wedge}$ are in the essential image. Therefore the functor $F$ is an equivalence of abelian categories. This shows condition (c). ∎
Corollary 3.6.
Let $(\mathcal{K},\mathcal{P},\mathcal{I})$ be a $d$-precluster tilting triple such that $\mathcal{B}^{\prime}=\tau_{\leq 0}\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ is abelian. Then the dg category $\mathcal{P}$ is concentrated in degree zero with weak kernels and the map sending $X\in\mathcal{B}^{\prime}$ to $\mathrm{Hom}_{\mathcal{B}^{\prime}}(-,X)|_{\mathcal{P}}$ extends to an exact equivalence $F:\mathcal{B}^{\prime}\xrightarrow{{}_{\sim}}\mathrm{mod}\mathcal{P}$.
Lemma 3.7.
Let $(\mathcal{K},\mathcal{P},\mathcal{I})$ be a $d$-precluster tilting triple such that $\mathcal{B}^{\prime}=\tau_{\leq 0}\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ is abelian. Then the map sending an object $X$ to $\mathrm{Hom}_{\mathcal{B}^{\prime}}(X,-)|_{\mathcal{I}}$ extends to a functor $G:\mathcal{B}^{\prime}\rightarrow(\mathcal{I}\mbox{-}\mathrm{mod})^{op}$, which is fully faithful when restricted to the full subcategory $H^{0}(\mathcal{K})$.
Proof.
Let $K$ and $K^{\prime}$ be objects in $\mathcal{K}$. Since $H^{0}(\mathcal{I})$ is covariantly finite in $H^{0}(\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d)})$, we have triangles in $\mathrm{tr}(\mathcal{K})$
$K\rightarrow I_{0}\rightarrow U\rightarrow\Sigma K,$
$U\rightarrow I_{1}\rightarrow V\rightarrow\Sigma U,$
where $I_{0}$ and $I_{1}$ are in $\mathcal{I}$ and $U$ and $V$ are in $\mathcal{B}^{\prime}$. Then we have short exact sequences of left $\mathcal{I}$-modules
$0\rightarrow\mathrm{Hom}(U,-)|_{\mathcal{I}}\rightarrow\mathrm{Hom}(I_{0},-)|_{\mathcal{I}}\rightarrow\mathrm{Hom}(K,-)|_{\mathcal{I}}\rightarrow 0,$
$0\rightarrow\mathrm{Hom}(V,-)|_{\mathcal{I}}\rightarrow\mathrm{Hom}(I_{1},-)|_{\mathcal{I}}\rightarrow\mathrm{Hom}(U,-)|_{\mathcal{I}}\rightarrow 0$
So we have an exact sequence in (-$(\mathcal{I}\mbox{-}\mathrm{mod})^{op}$
$0\rightarrow G(K)\rightarrow G(I_{0})\rightarrow G(I_{1})$
We also have exact sequences of Hom spaces in $\mathrm{tr}(\mathcal{K})$
$0=\mathrm{Hom}(K^{\prime},\Sigma^{-1}U)\rightarrow\mathrm{Hom}(K^{\prime},K)\rightarrow\mathrm{Hom}(K^{\prime},I_{0})\rightarrow\mathrm{Hom}(K^{\prime},U),$
$0=\mathrm{Hom}(K^{\prime},\Sigma^{-1}V)\rightarrow\mathrm{Hom}(K^{\prime},U)\rightarrow\mathrm{Hom}(K^{\prime},I_{1})\rightarrow\mathrm{Hom}(K^{\prime},V).$
So we have the following diagram
[…]
where the second row is a sequence of Hom spaces in (-$(\mathcal{I}\mbox{-}\mathrm{mod})^{op}$. So the functor $G|_{H^{0}(\mathcal{K})}$ is fully faithful.
3.8. Iyama–Solberg’s correspondence
In this subsection, we assume that $k$ is a commutative artin ring. We compare our correspondence with the correspondence obtained in [15, Theorem 4.5]. Let $\Lambda$ be an Artin $k$-algebra and $\mathcal{I}$ the full subcategory of injective $\Lambda$-modules.
In [14, 1.4.1], for $d\geq 1$, Iyama defined the functors
$\tau_{d}=\tau\Omega^{d-1}:\underline{\mathrm{mod}}\Lambda\rightarrow\overline{\mathrm{mod}}\Lambda\text{ and }\tau_{d}^{-}=\tau^{-}\Omega^{-(d-1)}:\overline{\mathrm{mod}}\Lambda\rightarrow\underline{\mathrm{mod}}\Lambda$
as the $d$-Auslander–Reiten translation.
Definition 3.9 ([15, Definition 3.2]).
A subcategory $\mathcal{C}$ of $\mathrm{mod}\Lambda$ is a $d$-precluster tilting subcategory if it satisfies the following conditions.
  • (i)
    $\mathcal{C}$ is a generator-cogenerator for $\mathrm{mod}\Lambda$.
  • (ii)
    $\tau_{d}(\mathcal{C})\subseteq\mathcal{C}$ and $\tau_{d}^{-}(\mathcal{C})\subseteq\mathcal{C}$.
  • (iii)
    $\mathrm{Ext}^{i}_{\Lambda}(\mathcal{C},\mathcal{C})=0$ for $0<i<d$.
  • (iv)
    $\mathcal{C}$ is a functorially finite subcategory of $\mathrm{mod}\Lambda$.
If moreover $\mathcal{C}$ admits an additive generator $M$, we say that $M$ is a $d$-precluster tilting module.
Put
${{}^{\perp_{d-1}}\mathcal{C}}=\{\text{$X$ in $\mathrm{mod}\Lambda\mid\mathrm{Ext}^{i}_{\Lambda}(X,\mathcal{C})=0$ for $0<i<d$}\},$
${\mathcal{C}^{\perp_{d-1}}}=\{\text{$X$ in $\mathrm{mod}\Lambda\mid\mathrm{Ext}^{i}_{\Lambda}(\mathcal{C},X)=0$ for $0<i<d$}\}.$
Suppose conditions (i), (iii) and (iv) hold and $d>1$, then by [15, Proposition 3.8 (b)], condition (ii) is equivalent to ${{}^{\perp_{d-1}}\mathcal{C}}={\mathcal{C}^{\perp_{d-1}}}$. Indeed, we have the following result. Here we include a proof for completeness.
Lemma 3.10.
Suppose $d>1$. Let $\mathcal{C}$ be an additive subcategory of $\mathrm{mod}\Lambda$ satisfying the conditions (i), (iii) and (iv) in Definition 3.9. Then $\tau_{d}(\mathcal{C})\subseteq\mathcal{C}$ (resp. $\tau_{d}^{-}(\mathcal{C})\subseteq\mathcal{C}$) if and only if ${{}^{\perp_{d-1}}\mathcal{C}}\subseteq\mathcal{C}^{\perp_{d-1}}$ (resp. ${{}^{\perp_{d-1}}\mathcal{C}}\supseteq\mathcal{C}^{\perp_{d-1}}$).
Proof.
We prove the case for $\tau_{d}$ and the case for $\tau_{d}^{-}$ is similar.
Assume $\tau_{d}(\mathcal{C})\subseteq\mathcal{C}$. Since $\Lambda$ is contained in $\mathcal{C}$, we have $\mathcal{C}\subseteq{{}^{\perp_{d-1}}\Lambda}$ by condition (iii). So we have
$\mathrm{Ext}^{i}_{\Lambda}(\mathcal{C},{{}^{\perp_{d-1}}\mathcal{C}})\xrightarrow{{}_{\sim}}D\mathrm{Ext}^{d-i}_{\Lambda}({{}^{\perp_{d-1}}}\mathcal{C},\tau_{d}(\mathcal{C}))=0$
for $0<i<d$. It follows that ${{}^{\perp_{d-1}}}\mathcal{C}\subseteq\mathcal{C}^{\perp_{d-1}}$.
Conversely, we assume that ${{}^{\perp_{d-1}}}\mathcal{C}\subseteq\mathcal{C}^{\perp_{d-1}}$. Since $D\Lambda$ is contained in $\mathcal{C}$, it follows that $\mathcal{C}^{\perp_{d-1}}\subseteq(D\Lambda)^{\perp_{d-1}}$. Hence ${{}^{\perp_{d-1}}}\mathcal{C}\subseteq(D\Lambda)^{\perp_{d-1}}$ and we have
$\mathrm{Ext}^{i}_{\Lambda}(\mathcal{C},{{}^{\perp_{d-1}}}\mathcal{C})\xrightarrow{{}_{\sim}}D\mathrm{Ext}^{d-i}_{\Lambda}(\tau_{d}^{-}({{}^{\perp_{d-1}}}\mathcal{C}),\mathcal{C})=0$
for $0<i<d$. Hence $\tau_{d}^{-}({{}^{\perp_{d-1}}}\mathcal{C})\subseteq{{}^{\perp_{d-1}}}\mathcal{C}$. Since ${{}^{\perp_{d-1}}}\mathcal{C}\subseteq(D\Lambda)^{\perp_{d-1}}$, we have
$\mathrm{Ext}^{i}_{\Lambda}(({{}^{\perp_{d-1}}}\mathcal{C})^{\perp_{d-1}},{{}^{\perp_{d-1}}}\mathcal{C})\xrightarrow{{}_{\sim}}D\mathrm{Ext}^{d-i}_{\Lambda}(\tau_{d}^{-}({{}^{\perp_{d-1}}}\mathcal{C}),({{}^{\perp_{d-1}}}\mathcal{C})^{\perp_{d-1}})=0$
for $0<i<d$. It follows that $({{}^{\perp_{d-1}}}\mathcal{C})^{\perp_{d-1}}\subseteq{{}^{\perp_{d-1}}({{}^{\perp_{d-1}}}\mathcal{C})}\subseteq{{}^{\perp_{d-1}}}\mathcal{C}$. We show that $({{}^{\perp_{d-1}}}\mathcal{C})^{\perp_{d-1}}=\mathcal{C}$. Let $X$ be an object in $({{}^{\perp_{d-1}}}\mathcal{C})^{\perp_{d-1}}$. By conditions (i) and (iv), we have an exact sequence
[…]
where $f$ is a left $\mathcal{C}$-approximation of $X$. Then $Y\in{{}^{\perp_{d-1}}}\mathcal{C}$ and the sequence (3.10.1) splits. Hence we have $X\in\mathcal{C}$ and $({{}^{\perp_{d-1}}}\mathcal{C})^{\perp_{d-1}}=\mathcal{C}$. Since $\mathcal{C}\subseteq{{}^{\perp_{d-1}}\Lambda}$, by the assumption that ${{}^{\perp_{d-1}}}\mathcal{C}\subseteq\mathcal{C}^{\perp_{d-1}}$ we have
$\mathrm{Ext}^{i}_{\Lambda}({{}^{\perp_{d-1}}}\mathcal{C},\tau_{d}(\mathcal{C}))\xrightarrow{{}_{\sim}}D\mathrm{Ext}^{d-i}_{\Lambda}(\mathcal{C},{{}^{\perp_{d-1}}}\mathcal{C})=0$
for $0<i<d$. It follows that $\tau_{d}(\mathcal{C})\subseteq({{}^{\perp_{d-1}}}\mathcal{C})^{\perp_{d-1}}=\mathcal{C}$ and we are done. ∎
Let $(\mathcal{K},\mathcal{P},\mathcal{I})$ be a $d$-precluster tilting triple which satisfies conditions (a), (b) and (c) in Proposition 3.5. By Corollary 3.6, the connective exact dg category $\tau_{\leq 0}\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ can be identified with the abelian category $\mathrm{mod}\mathcal{P}$. Then the category $\mathcal{I}$ can be identified with the full subcategory of projective-injectives in $\mathrm{mod}\mathcal{P}$, and the dg category $\mathcal{K}$ can be identified with the full dg subcategory of $\mathcal{D}^{b}_{dg}(\mathrm{mod}\mathcal{P})$ consisting of all $(d+1)$-st syzygies of $\mathcal{P}$-modules. So $\mathcal{K}$ and $\mathcal{I}$ are determined by $\mathcal{P}$ if $\mathcal{P}$ is part of a $d$-precluster tilting triple $(\mathcal{K},\mathcal{P},\mathcal{I})$ such that $\mathcal{B}^{\prime}=\tau_{\leq 0}\mathcal{B}_{\mathcal{K},\mathcal{P},\mathcal{I}}^{(d+1)}$ is abelian.
Now, suppose that $\mathcal{P}=\mathrm{add}M$ for some $\Lambda$-module $M$ where $\Lambda$ is an Artin algebra, and by Lemma 3.7, we also assume that $\mathcal{I}=\mathrm{inj}\Lambda$. We identify $\mathcal{P}$, respectively $\mathcal{I}$, with their essential images under the Yoneda embedding $\mathcal{P}\hookrightarrow\mathrm{mod}\mathcal{P}$. We further identify them with their essential images under the dg functor $\mathrm{mod}\mathcal{P}\rightarrow\mathcal{D}^{b}_{dg}(\mathrm{mod}\mathcal{P})$ and view them as full dg subcategories of $\mathcal{K}$, which is a full dg subcategory of $\mathcal{D}^{b}_{dg}(\mathrm{mod}\mathcal{P})$ consisting of all $(d+1)$-st syzygies of $\mathcal{P}$-modules. We claim that the triple $(\mathcal{K},\mathcal{P},\mathcal{I})$ is $d$-precluster tilting satisfying the conditions in Proposition 3.5 if and only if the $\Lambda$-module $M$ is $d$-precluster tilting. We identify $\mathrm{pretr}(\mathcal{K})$ with $\mathcal{D}^{b}_{dg}(\mathrm{mod}\mathcal{P})$ via the canonical quasi-equivalence $\mathrm{pretr}(\mathcal{K})\xrightarrow{{}_{\sim}}\mathcal{D}^{b}_{dg}(\mathrm{mod}\mathcal{P})$ induced by the inclusion $\mathcal{K}\hookrightarrow\mathcal{D}^{b}_{dg}(\mathrm{mod}\mathcal{P})$.
Lemma 3.11.
Assume that $M$ is a $d$-precluster tilting $\Lambda$-module. Then the triple $(\mathcal{K},\mathcal{P},\mathcal{I})$ defined above is $d$-precluster tilting and satisfies conditions (a), (b) and (c) in Proposition 3.5.
Proof.
Property (1) in Definition 2.4 and conditions (a) and (c) and the “if” part of condition (b) are straightforward to check. Property (5) follows from the definition of $\mathcal{K}$ and the fact that a short exact sequence of $\mathcal{P}$-modules gives rise to a triangle in $\mathcal{D}^{b}(\mathrm{mod}\mathcal{P})\xrightarrow{{}_{\sim}}\mathrm{tr}(\mathcal{K})$. Property (2) is immediate because $\mathcal{I}=\mathrm{inj}\Lambda$ has an additive generator.
It remains to show properties (3) and (4) and the “only if” part of condition (b). We show them in the $d=1$ case and $d>1$ case separately. Since $\mathcal{P}$ is contravariantly finite in $\mathrm{mod}\Lambda$ and contains $\Lambda$, the functor
$F:\mathrm{mod}\Lambda\rightarrow\mathrm{mod}\mathcal{P},\quad X\mapsto\mathrm{Hom}_{\mathrm{mod}\Lambda}(-,X)|_{\mathcal{P}}$
is fully faithful. Each object $K$ in $\mathcal{K}$ is of the form $F(K_{d+1})$ for some complex in $\mathrm{mod}\Lambda$
[…]
where $P_{i}\in\mathcal{P}$ for $0\leq i\leq d$, which remains exact after applying the functor $\mathrm{Hom}_{\mathrm{mod}\Lambda}(M,-)$.
Suppose $d=1$. Then the objects in $\mathcal{K}$ are precisely the $\mathcal{P}$-modules $\mathrm{Hom}_{\mathrm{mod}\Lambda}(-,K)|_{\mathcal{P}}$ where $K\in\mathrm{mod}\Lambda$: each $\Lambda$-module $K$ can be realized as the kernel of a morphism between objects in $\mathcal{P}$ since $\mathcal{P}$ contains an injective cogenerator. The “only if” part of condition (b) is straightforward to check. Let $K$ be any object in $\mathrm{mod}\Lambda$. By items (i) and (iv), we have an exact sequence in $\mathrm{mod}\Lambda$
[…]
where $P\in\mathcal{P}=\mathrm{add}M$, which remains exact after applying the functor $\mathrm{Hom}_{\mathrm{mod}\Lambda}(M,-)$. So any morphism $K^{\prime}\rightarrow\tau M$ factors through the morphism $K^{\prime}\rightarrow P$, cf. [2, Chapter III, Corollary 4.2]. Since $\tau(\mathcal{P})\subseteq\mathcal{P}$ and $\tau^{-}(\mathcal{P})\subseteq\mathcal{P}$, the functor $\tau:\underline{\mathcal{P}}\rightarrow\overline{\mathcal{P}}$ is an equivalence of categories. It follows that any morphism $K^{\prime}\rightarrow M$ factors through the morphism $K^{\prime}\rightarrow P$. Therefore we have $\mathrm{Ext}^{\geq 1}_{\mathrm{mod}\mathcal{P}}(\mathrm{Hom}_{\mathrm{mod}\Lambda}(-,K)|_{\mathcal{P}},\mathrm{Hom}_{\mathrm{mod}\Lambda}(-,M))=0$. This proves property (4). Property (3) also follows by applying the Horseshoe Lemma.
Suppose $d>1$. Then we have ${{}^{\perp_{d-1}}\mathcal{P}}={\mathcal{P}^{\perp_{d-1}}}$. Consider a complex (3.5.1) in $H^{0}(\mathcal{K})$ such that the corresponding complex (3.5.3) is exact for each $P\in\mathcal{P}$. Since $G$ is fully faithful, it is given by a complex (3.11.1) under the functor $G$ where $P_{i}\in\mathcal{P}$ for $0\leq i\leq d$, which remains exact after applying the functor $\mathrm{Hom}_{\mathrm{mod}\Lambda}(M,-)$. Since $\Lambda$ is contained in $\mathcal{P}$, the complex (3.11.1) is exact in $\mathrm{mod}\Lambda$. Since $\mathcal{I}$ consists of injective $\Lambda$-modules and $G$ is fully faithful, we see that the corresponding complex (3.5.2) is exact for each $I\in\mathcal{I}$. This shows the “only if” part of condition (b). Let us show property (4). Let $K$ be an object in $\mathcal{K}$. So it is of the form $F(K_{d+1})$ for some complex (3.11.1) in $\mathrm{mod}\Lambda$ where $P_{i}\in\mathcal{P}$ for $0\leq i\leq d$, which remains exact after applying the functor $\mathrm{Hom}_{\mathrm{mod}\Lambda}(M,-)$. Then we have $K_{d+1}\in{\mathcal{P}^{\perp_{d-1}}}$. Since we have ${{}^{\perp_{d-1}}\mathcal{P}}={\mathcal{P}^{\perp_{d-1}}}$, the object $K_{d+1}$ lies in ${{}^{\perp_{d-1}}\mathcal{P}}$. Since $\mathcal{P}=\mathrm{add}M$ is contravariantly finite in $\mathrm{mod}\Lambda$, we have an exact sequence in $\mathrm{mod}\Lambda$
$\cdots\rightarrow P_{d+3}\xrightarrow{g_{d+2}}P_{d+2}\xrightarrow{g_{d+1}}P_{d+1}\xrightarrow{g_{d}}K_{d+1}\rightarrow 0$
which is exact after applying the functor $G$. Similarly we have that $\mathrm{Im}(g_{i})\in{{}^{\perp_{d-1}}\mathcal{P}}$ for $i\geq d+1$. It follows that for each $P\in\mathcal{P}$ we have
$\mathrm{Ext}^{\geq 1}_{\mathrm{mod}\mathcal{P}}(\mathrm{Hom}_{\mathrm{mod}\Lambda}(-,K_{d+1})|_{\mathcal{P}},\mathrm{Hom}_{\mathrm{mod}\Lambda}(-,P))=0.$
This shows property (4). By condition (iv), we have a left $\mathcal{P}$-approximation $K_{d+1}\rightarrowtail Q_{d}$ with $L_{d}$ its cokernel. Again we have a left $\mathcal{P}$-approximation $L_{d}\rightarrowtail Q_{d-1}$. We repeat the above procedure and we have a complex in $\mathrm{mod}\Lambda$
[…]
where $Q_{i}\in\mathcal{P}$ for $0\leq i\leq d$. Since $K_{d+1}\in{\mathcal{P}^{\perp_{d-1}}}$, the above complex remains exact after applying the functor $\mathrm{Hom}(M,-)$. So we may replace the complex (3.11.1) by (3.11.2). Since $f_{d+1}$ is a left $\mathcal{P}$-approximation, we have
$\mathrm{Ext}^{i}_{\mathrm{mod}\mathcal{P}}(\mathrm{Hom}_{\mathrm{mod}\Lambda}(-,L_{d})|_{\mathcal{P}},\mathrm{Hom}_{\mathrm{mod}\Lambda}(-,M))=0.$
for $11\leq i\leq d-$ 1. So we may apply the Horseshoe Lemma and the dg category $\mathcal{K}$ is extension-closed in $\mathrm{pretr}(\mathcal{K})\xrightarrow{{}_{\sim}}\mathcal{D}^{b}_{dg}(\mathrm{mod}\mathcal{P})$. This shows property (3) and hence finishes the proof. ∎
Lemma 3.12.
Assume the triple $(\mathcal{K},\mathcal{P},\mathcal{I})$ is $d$-precluster tilting and satisfies the conditions (a), (b) and (c) in Proposition 3.5. Then the $\Lambda$-module $M$ is $d$-precluster tilting.
Proof.
Item (iv) in Definition 3.9 is straightforward. Since $\mathcal{P}=\mathrm{add}M$ contains the injective $\Lambda$-modules, it is a cogenerator in $\mathrm{mod}\Lambda$. Let $I$ be an object in $\mathcal{I}$. We show that $I^{\wedge}$ is injective in $\mathrm{mod}\mathcal{P}$. Let $X$ be an object in $\mathrm{mod}\mathcal{P}$. By condition (c) in Proposition 3.5 (and condition (a) if $d=1$), we have an exact sequence in $\mathrm{mod}\mathcal{P}$
[…]
By applying the functor $\mathrm{Hom}_{\mathrm{mod}\mathcal{P}}(-,I^{\wedge})$, we get a sequence of Hom spaces in $\mathrm{mod}\mathcal{P}$
[…]
which is exact by condition (b). It follows that $\mathrm{Ext}^{1}_{\mathrm{mod}\mathcal{P}}(X,I^{\wedge})=0$ and hence $I^{\wedge}$ is injective in $\mathrm{mod}\mathcal{P}$. Let us show that $\mathcal{P}$ also contains the projective $\Lambda$-modules. The canonical morphism $\mathrm{Hom}_{\mathrm{mod}\Lambda}(\Lambda,I)\rightarrow\mathrm{Hom}_{\mathrm{mod}\mathcal{P}}(\mathrm{Hom}_{\mathrm{mod}\Lambda}(-,\Lambda)|_{\mathcal{P}},I^{\wedge})$ is an isomorphism for $I\in\mathcal{I}$. Indeed, since $I^{\wedge}$ is injective, it is isomorphic to $D\mathrm{Hom}_{\mathrm{mod}\Lambda}(P,-)$ for some $P\in\mathcal{P}$. Then we have
$\mathrm{Hom}_{\mathrm{mod}\mathcal{P}}(\mathrm{Hom}_{\mathrm{mod}\Lambda}(-,\Lambda)|_{\mathcal{P}},I^{\wedge})$ $\xrightarrow{{}_{\sim}}\mathrm{Hom}_{\mathrm{mod}\mathcal{P}}(\mathrm{Hom}_{\mathrm{mod}\Lambda}(-,\Lambda)|_{\mathcal{P}},D\mathrm{Hom}(P,-))$
$\xrightarrow{{}_{\sim}}D\mathrm{Hom}(P,\Lambda)$
$\xrightarrow{{}_{\sim}}\mathrm{Hom}(\Lambda,I).$
We have a sequence in $\mathrm{mod}\Lambda$
[…]
where $Q_{1}$ and $Q_{0}$ are in $\mathcal{P}$, which becomes exact after applying the functor $\mathrm{Hom}_{\mathrm{mod}\Lambda}(M,-)$. Since $I^{\wedge}$ is injective in $\mathrm{mod}\Lambda$ for each injective $I\in\mathrm{mod}\Lambda$, by the above we have an exact sequence
[…]
It follows that the sequence (3.12.1) is exact in $\mathrm{mod}\Lambda$ and hence $\Lambda$ is a direct summand of $Q_{0}$. Therefore we have $\Lambda\in\mathcal{P}$ and this proves item (i) in Definition 3.9. As a consequence, the functor
$G:\mathrm{mod}\Lambda\rightarrow\mathrm{mod}\mathcal{P},\quad X\mapsto\mathrm{Hom}_{\mathrm{mod}\Lambda}(-,X)|_{\mathcal{P}}$
is fully faithful.
Let us show item (iii). We take the first $d$-term of an injective coresolution of $M$ in $\mathrm{mod}\Lambda$
[…]
By condition (b) in Proposition 3.5, the sequence (3.12.2) remains exact after applying the functor $\mathrm{Hom}_{\mathrm{mod}\Lambda}(M,-)$. So by definition we have $\mathrm{Ext}^{i}(M,M)=0$ for $0<i<d$.
Finally let us show item (ii). We will show this in $d=1$ case and $d>1$ case separately. Assume $d=1$. Let us show that $\tau^{-}(\mathcal{P})\subset\mathcal{P}$. Since $\mathcal{P}=\mathrm{add}M$ contains the injectives and the projectives, and the number of isoclasses of indecomposable injectives equals the number of isoclasses of indecomposable projectives, it follows that $\tau(\mathcal{P})\subseteq\mathcal{P}$. Let $N$ be an indecomposable noninjective $\Lambda$-module in $\mathcal{P}$. We have $\mathrm{Ext}^{\geq 1}_{\mathrm{mod}\mathcal{P}}(G(\tau^{-1}(N)),G(P))=0$ for $P\in\mathcal{P}$ by property (4) in Definition 2.4. We have a sequence in $\mathrm{mod}\Lambda$
[…]
where $P^{i}\in\mathcal{P}$ for $i\geq 0$, and which is exact after applying the functor $G$. So after applying the functor $G$, it is a projective resolution of $G(\tau^{-1}(N))$. It follows that the above sequence (3.12.3) is exact after applying the functor $\mathrm{Hom}_{\mathrm{mod}\Lambda}(-,P)$ for each $P\in\mathcal{P}$. Put $L=\mathrm{Im}(u)$. Then the inclusion of $L$ into $P^{0}$ induces an epimorphism $\mathrm{Hom}(P^{0},P)\rightarrow\mathrm{Hom}(L,P)$ for each $P\in\mathcal{P}$. Hence the morphism $v:P^{0}\rightarrow\tau^{-}(N)$ induces an epimorphism $\mathrm{Hom}(\tau^{-}(P),P^{0})\rightarrow\mathrm{Hom}(\tau^{-}(P),\tau^{-}(N))$. In particular, the morphism $v$ splits and $\tau^{-}(N)$ is a direct summand of $P^{0}\in\mathcal{P}$.
Assume $d>1$. Let us show that $\tau_{d}^{-}(\mathcal{P})\subseteq\mathcal{P}$. By [14, Theorem 1.4.1], it follows that $\tau_{d}(\mathcal{P})\subseteq\mathcal{P}$. By Lemma 3.10, it suffices to show ${{}^{\perp_{d-1}}\mathcal{P}}\supseteq\mathcal{P}^{\perp_{d-1}}$. Let $Y\in\mathcal{P}^{\perp_{d-1}}$. Then we have an exact sequence in $\mathrm{mod}\Lambda$
$0\rightarrow Y\rightarrow Q_{d}\rightarrow Q_{d-1}\rightarrow\cdots\rightarrow\mathcal{Q}_{1}\rightarrow Q_{0},$
where $Q_{d}\in\mathcal{P}$, and which is exact after applying the functor $\mathrm{Hom}(-,M)$. Since $Y\in\mathcal{P}^{\perp_{d-1}}$, the above sequence remains exact after applying the functor $\mathrm{Hom}(P,-)$ for each $P\in\mathcal{P}$. Hence $G(Y)$ lies in $\mathcal{K}$. By property (4) of Definition 2.4, we have $\mathrm{Ext}^{\geq 1}_{\mathrm{mod}\mathcal{P}}(G(Y),G(P))=0$ for each $P\in\mathcal{P}$. We also have an exact sequence in $\mathrm{mod}\Lambda$
$\cdots\rightarrow Q_{d+3}\xrightarrow{g_{d+2}}Q_{d+2}\xrightarrow{g_{d+1}}Q_{d+1}\xrightarrow{g_{d}}Y\rightarrow 0$
where $Q_{i}\in\mathcal{P}$ for $i\geq d+1$, and which remains exact after applying the functor $\mathrm{Hom}(P,-)$ for each $P\in\mathcal{P}$. So we get a projective resolution of $G(Y)$ after applying the functor $G$. Hence the above sequence remains exact after applying the functor $\mathrm{Hom}(-,P)$ for each $P\in\mathcal{P}$. Since $\mathrm{Ext}^{i}_{\Lambda}(\mathcal{P},\mathcal{P})=0$ for $0<i<d$, it follows that $Y\in{{}^{\perp_{d-1}}\mathcal{P}}$. ∎
3.13. Connection to higher Auslander–Solberg correspondence for exact categories
In the rest of this section, we assume $d\geq 1$ and compare our correspondence with the correspondence obtained by Grevstad in [11]. By combining Proposition 3.18 and Lemma 3.19, we see that we get a generalization of the correspondence [11, Theorem 3.8] in the case of weakly idempotent complete exact categories with enough injectives.
Definition 3.14 ([7, Proposition 1.7]).
Let $\mathcal{M}\subseteq\mathcal{E}$ be a full subcategory of an exact category. The relative exact structure $F_{\mathcal{M}}$ is given by the conflations in $\mathcal{E}$ which remain exact after applying the functor $\mathrm{Hom}_{\mathcal{E}}(M,-)$ for any $M\in\mathcal{M}$.
Definition 3.15 ([11]).
Let $\mathcal{M}$ be a subcategory of an exact category $\mathcal{E}$. We say that $\mathcal{M}$ is partial $d$-precluster tilting if it satisfies
  • (a)
    every object $E\in\mathcal{E}$ has a left $\mathcal{M}$-approximation by an inflation $E\rightarrowtail M$;
  • (b)
    every object $E\in\mathcal{E}$ has a right $\mathcal{M}$-approximation by a deflation $M^{\prime}\twoheadrightarrow E$
  • (c)
    $\mathcal{M}$ is $d$-rigid, i.e. $\mathrm{Ext}^{1\sim d-1}_{\mathcal{E}}(\mathcal{M},\mathcal{M})=0$;
  • (d)
    the relative injective dimension $\mathrm{id}_{F_{\mathcal{M}}}M$ is strictly less than $d$ for any $M\in\mathcal{M}$, where $F_{\mathcal{M}}$ is the exact substructure defined in Definition 3.14;
A morphism $f:X\rightarrow Y$ in an exact category $\mathcal{E}$ is admissible if $f$ admits a factorization $X\twoheadrightarrow I\rightarrowtail Y$. A cochain complex over $\mathcal{E}$
$\ldots\rightarrow X^{n-1}\xrightarrow{d^{n-1}}X^{n}\xrightarrow{d^{n}}X^{n+1}\rightarrow\ldots$
is called acyclic or exact if each $d^{i}$ is admissible and $\mathrm{ker}(d^{i+1})=\mathrm{im}(d^{i})$. Let $\mathcal{M}$ be a full subcategory of $\mathcal{E}$. We write $\mathrm{mod}_{adm}(\mathcal{M})$ [9, Definition 3.1 (2)] for the full subcategory of $\mathrm{Mod}\,\mathcal{M}$ consisting of those functors $F$ that admit a projective presentation
[…]
for some morphism $f:M\rightarrow N$ in $\mathcal{M}$ which is admissible in $\mathcal{E}$. Recall that for a full subcategory $\mathcal{X}$ of $\mathcal{E}$, we write
${{}^{\perp}\mathcal{X}}=\{E\in\mathcal{E}\mid\mathrm{Hom}_{\mathcal{E}}(E,X)=0\text{ for any $X\in\mathcal{X}$}\}.$
Put $\mathrm{cogen}(\mathcal{X})=\{E\in\mathcal{E}\mid\exists\text{ inflation }E\rightarrowtail X\text{ with }X\in\mathcal{X}\}$. A pair $(\mathcal{T},\mathcal{F})$ of full subcategories of $\mathcal{E}$ is a torsion pair [12, Definition 2.17] if
  • (1)
    For every $T\in\mathcal{T}$ and $F\in\mathcal{F}$, we have $\mathrm{Hom}_{\mathcal{E}}(T,F)=0$.
  • (2)
    For any object $M\in\mathcal{E}$, there exists a conflation $T\rightarrowtail M\twoheadrightarrow F$ with $T\in\mathcal{T}$ and $F\in\mathcal{F}$.
Theorem 3.16 ([11]).
There is a bijection between the following:
  • (1)
    Equivalence classes of partial $d$-precluster tilting subcategories $\mathcal{M}$ of exact categories $\mathcal{E}$ with enough injectives.
  • (2)
    Equivalence classes of exact categories $\mathcal{E}^{\prime}$ with enough projectives, where we denote by $\mathcal{P}=\mathrm{proj}\,(\mathcal{E}^{\prime})$ the full subcategory of projectives, satisfying the following conditions
    ($a$) $\mathrm{domdim}(\mathcal{E}^{\prime})\geq d+1$.
    ($b$) Each projective $P$ has injective dimension $\leq d+1$, i.e. $\mathrm{Ext}^{\geq d+2}_{\mathcal{E}^{\prime}}(-,P)=0$.
    ($c$) Any morphism $X\rightarrow E$ with $E\in{{}^{\perp}\mathcal{P}}$ is admissible.
    ($d$) $\mathcal{P}$ is admissibly covariantly finite, i.e. for each object $E$ in $\mathcal{E}$, there is a left $\mathcal{P}$-approximation $E\rightarrow P$ which is admissible.
The bijection from (1) to (2) sends $(\mathcal{E},\mathcal{M})$ to $\mathrm{mod}_{adm}(\mathcal{M})$.
Remark 3.17.
The above bijection is a restriction of the bijection in [11, Theorem 3.8]. More precisely, in Proposition 6.3 in loc. cit, it is shown that in the bijection of Theorem 3.8 in loc. cit, the exact categories $\mathcal{E}$ has enough injectives if and only if for the corresponding exact category $\mathcal{E}^{\prime}$ we have $\mathrm{domdim}(\mathcal{E}^{\prime})\geq 1$. But since $\mathcal{M}$ is $d$-rigid, we have $\mathrm{domdim}(\mathcal{E}^{\prime})\geq d+1$ in this case. By [12, Lemma 4.17 (2)], if $\mathcal{E}^{\prime}$ has enough projectives and if $\mathcal{P}$ is admissibly covariantly finite and $\mathrm{domdim}(\mathcal{E}^{\prime})\geq 1$, then $({{}^{\perp}\mathcal{P}},\mathrm{cogen}(\mathcal{P}))$ is a torsion pair. By [12, Lemma 4.15], if $\mathcal{E}^{\prime}$ has enough projectives and $\mathrm{domdim}(\mathcal{E}^{\prime})\geq 1$, then ${{}^{\perp}\mathcal{P}}$ is closed under admissible subobjects. Since $\mathrm{domdim}(\mathcal{E}^{\prime})\geq d+1$, a similar proof of [12, Lemma 4.14] shows that $\mathrm{Ext}^{1\sim d}_{\mathcal{E}}({{}^{\perp}\mathcal{P}},\mathcal{P})=0$. Therefore the exact categories appearing in (2) of the above bijection are partial $d$-minimal Auslander–Gorenstein in the sense of [11, Definition 3.4].
Let $\mathcal{E}$ be a weakly idempotent complete exact category with enough injectives. Let $\mathcal{M}$ be a $d$-precluster tilting subcategory and $\mathcal{I}$ the full subcategory of $\mathcal{E}$ on the injectives. We define $\mathcal{K}$ to be the full dg subcategory $\mathcal{D}^{b}_{dg}(\mathrm{Mod}\,\mathcal{M})$ consisting of the right $\mathcal{M}$-modules $\mathrm{Hom}_{\mathcal{E}}(-,K)|_{\mathcal{M}}$ where $K$ appears in an exact sequence
[…]
where $M_{i}\in\mathcal{M}$ for $0\leq i\leq d$ and which remains exact after applying the functor $\mathrm{Hom}_{\mathcal{E}}(M,-)$ for each $M\in\mathcal{M}$. We identify $\mathcal{M}$ with the full subcategory of representable $\mathcal{M}$-modules. By Proposition 3.18 and Lemma 3.19 below, we have the following diagram
[…]
Proposition 3.18.
The triple $(\mathcal{K},\mathcal{M},\mathcal{I})$ is $d$-precluster tilting and satisfies the conditions in Proposition 3.2.
Proof.
The inclusion $\mathcal{K}\hookrightarrow\mathcal{D}^{b}_{dg}(\mathrm{Mod}\,\mathcal{M})$ induces an inclusion $\mathrm{pretr}(\mathcal{K})\hookrightarrow\mathcal{D}^{b}_{dg}(\mathrm{Mod}\,\mathcal{M})$. It is clear that $\mathcal{K}$ is concentrated in non-negative degrees. So condition (a) in Proposition 3.2 and property (1) in Definition 2.4 hold.
By Definition 3.15 (b), the functor $F:\mathcal{E}\rightarrow\mathrm{Mod}\,\mathcal{M}$ sending $X\in\mathcal{E}$ to $\mathrm{Hom}_{\mathcal{E}}(-,X)|_{\mathcal{M}}$ is fully faithful. We thus identify $H^{0}(\mathcal{K})$ with the corresponding full subcategory of $\mathcal{E}$. By definition of $\mathcal{K}$ and that $\mathcal{M}$ is $d$-rigid, we have that $\mathrm{Ext}^{1\sim d-1}_{\mathcal{E}}(M,K)=0$ for $M\in\mathcal{M}$ and $K\in H^{0}(\mathcal{K})$. Let us show that $\mathrm{Ext}^{1}_{F_{\mathcal{M}}}(K,M)=0$ for $K\in H^{0}(\mathcal{K})$ and $M\in\mathcal{M}$. Indeed, we have conflations in $F_{\mathcal{M}}$
[…]
$\ldots$
[…]
where $N_{i}\in\mathcal{M}$ for $2\leq i\leq d$. Then we have
$\mathrm{Ext}^{1}_{F_{\mathcal{M}}}(K,M)\xrightarrow{{}_{\sim}}\mathrm{Ext}^{2}_{F_{\mathcal{M}}}(V_{d},M)\xrightarrow{{}_{\sim}}\ldots\xrightarrow{{}_{\sim}}\mathrm{Ext}^{d}_{F_{\mathcal{M}}}(V_{2},M)=0$
since the relative injective dimension $\mathrm{id}_{F_{\mathcal{M}}}M$ is strictly less than $d$ for $M\in\mathcal{M}$.
Now we show that $\mathrm{Ext}^{\geq 1}_{\mathrm{Mod}\,\mathcal{M}}(F(K),F(M))=0$ for $K\in H^{0}(\mathcal{K})$ and $M\in\mathcal{M}$ and so property (4) in Definition 2.4 holds. By Definition 3.15 (b), there exists a conflation in $\mathcal{E}$
[…]
where $p$ is a right $\mathcal{M}$-approximation of $K$. It follows that $K^{\prime}$ also lies in $H^{0}(\mathcal{K})$. Therefore property (5) in Definition 2.4 holds. Now the conflation (3.18.1) also lies in $F_{\mathcal{M}}$. Since $\mathrm{Ext}^{1}_{F_{\mathcal{M}}}(K,M)=0$, the morphism $i$ induces a surjection $\mathrm{Hom}_{\mathcal{E}}(M^{\prime},M)\rightarrow\mathrm{Hom}_{\mathcal{E}}(K^{\prime},M)$. So we have an exact sequence in $F_{\mathcal{M}}$
[…]
which remains exact after applying the functor $\mathrm{Hom}_{\mathcal{E}}(-,M)$ for each $M\in\mathcal{M}$. After applying the functor $F:\mathcal{E}\rightarrow\mathrm{Mod}\,\mathcal{M}$, we obtain a projective resolution of $F(K)$. Since the functor $F$ is fully faithful when restricted to $H^{0}(\mathcal{K})$, we conclude that $\mathrm{Ext}^{\geq 1}_{\mathrm{Mod}\,\mathcal{M}}(F(K),F(M))=0$.
Condition (b) in Proposition 3.2 can be checked as follows: consider a complex
[…]
where $K_{d+1}\in H^{0}(\mathcal{K})$ and $M_{i}\in\mathcal{M}$ for $0\leq i\leq d$. Suppose that it remains exact after applying the functor $\mathrm{Hom}_{\mathcal{E}}(-,I)$ for each $I\in\mathcal{I}$. Since $\mathcal{E}$ has enough injectives and is weakly idempotent complete, we see that the morphism $K_{d+1}\rightarrow M_{d}$ is an inflation which is part of a conflation in $\mathcal{E}$
[…]
Then the map $M_{d}\rightarrow M_{d-1}$ factors through $U_{d}$ uniquely and similarly the map $U_{d}\rightarrow M_{d-1}$ is an inflation. We repeat the above argument and we see that the above sequence (3.18.2) is exact. Since $\mathcal{M}$ is $d$-rigid and $\mathrm{Ext}^{1\sim d-1}_{\mathcal{E}}(M,K_{d+1})=0$ for $M\in\mathcal{M}$, we see that the sequence (3.18.2) remains exact after applying the functor $\mathrm{Hom}_{\mathcal{E}}(M,-)$ for each $M\in\mathcal{M}$.
Let $K$ be any object in $H^{0}(\mathcal{K})$. Since $\mathcal{E}$ has enough injectives, we have an exact sequence
[…]
where $I_{i}\in\mathcal{I}$ for $0\leq i\leq d$. By the above, we see that the above sequence remains exact after applying the functor $\mathrm{Hom}(M,-)$ for each $M\in\mathcal{M}$. Hence, in the sequence (3.17.1), we could assume that $M_{i}\in\mathcal{I}$ for $0\leq i\leq d$. Therefore, by the Horseshoe Lemma we see that $\mathcal{K}$ is extension-closed in $\mathrm{pretr}(\mathcal{K})\subseteq\mathcal{D}^{b}_{dg}(\mathrm{Mod}\,\mathcal{M})$ and so property (3) in Definition 2.4 holds.
Let us show property (2), i.e. $H^{0}(\mathcal{I})$ is covariantly finite in
$\mathcal{B}_{\mathcal{K},\mathcal{M},\mathcal{I}}^{(d)}=\mathcal{K}\ast\Sigma\mathcal{K}\ast\ldots\ast\Sigma^{d}\mathcal{K}=\mathcal{M}\ast\Sigma\mathcal{M}\ast\ldots\ast\Sigma^{d-1}\mathcal{M}\ast\Sigma^{d}\mathcal{K}.$
Objects in $\mathcal{B}_{\mathcal{K},\mathcal{M},\mathcal{I}}^{(d)}$ are given by exact sequences in $\mathcal{E}$
[…]
where $K\in H^{0}(\mathcal{K})$ and $M_{i}\in\mathcal{M}$ for $0\leq i\leq d$. Indeed, the above sequence splits into conflations
[…]
$\ldots$
[…]
And the object in $\mathcal{B}_{\mathcal{K},\mathcal{M},\mathcal{I}}^{(d)}$, which corresponds to the exact sequence (3.18.3), is given by the cokernel of the morphism $F(U_{2})\rightarrow F(M_{1})$. Since the morphism $f_{2}:M_{2}\rightarrow M_{1}$ is admissible, we take its cokernel $M_{1}\rightarrow L$ and we have an inflation $L\rightarrowtail I$ for some $I\in\mathcal{I}$. Then the composition $M_{1}\rightarrow I$ gives rise to a morphism from the sequence (3.18.3) to $I$. It is straightforward to check that it is a left $\mathcal{I}$-approximation. ∎
Lemma 3.19.
As above, the functor $F:\mathcal{B}^{\prime}=\tau_{\leq 0}\mathcal{B}_{\mathcal{K},\mathcal{M},\mathcal{I}}^{(d+1)}\rightarrow\mathrm{Mod}\,\mathcal{M}$ in Lemma 3.3 induces an equivalence of exact categories $F:\mathcal{B}^{\prime}\rightarrow\mathrm{mod}_{adm}(\mathcal{M})$.
Proof.
Recall that we identify $H^{0}(\mathcal{K})$ with the corresponding subcategory of $\mathcal{E}$ via the functor $\mathcal{E}\rightarrow\mathrm{Mod}\,\mathcal{M}$ given by sending $E\in\mathcal{E}$ to $\mathrm{Hom}_{\mathcal{E}}(-,E)|_{\mathcal{M}}$. Let $X$ be an object in $\mathcal{B}^{\prime}$ with conflations (3.2.4) where $K_{d+1}\in H^{0}(\mathcal{K})$ and $P_{i}\in\mathcal{M}$ for $0\leq i\leq d$. Then the object $F(X)\in\mathrm{Mod}\,\mathcal{P}$ is given by the cokernel of the morphism $P_{1}^{\wedge}\rightarrow P_{0}^{\wedge}$. It suffices to show that the morphism $P_{1}\rightarrow P_{0}$ is admissible in $\mathcal{E}$. We can check this in a same way as we check condition (b) in the proof of Proposition 3.18. Therefore the functor $F$ gives rise to an exact functor $F:\mathcal{B}^{\prime}\rightarrow\mathrm{mod}_{adm}(\mathcal{M})$. Conversely, for each object $U$ in $\mathrm{mod}_{adm}(\mathcal{M})$, it is the cokernel of $P_{1}^{\wedge}\xrightarrow{f^{\wedge}}P_{0}^{\wedge}$ for some admissible morphism $f:P_{1}\rightarrow P_{0}$ in $\mathcal{M}$. By Definition 3.15 (2), we may complete it to an exact sequence in $\mathcal{E}$
[…]
where $P_{i}\in\mathcal{M}$ for $0\leq i\leq d$. It follows that $K_{d+1}\in H^{0}(\mathcal{K})$ and the exact sequence (LABEL:exact:K) gives rise to conflations (3.2.4) in $\mathcal{B}^{\prime}$. By definition, we have that $F(X)$ is isomorphic to the cokernel of $P_{1}^{\wedge}\xrightarrow{f^{\wedge}}P_{0}^{\wedge}$, which is in turn isomorphic to $U$. ∎